### v2

# Differential graded orders, their class groups and idèles

Abstract.
For a Dedekind domain $R$ with field of fractions $K$ a classical $R$-order in a semisimple $K$-algebra $A$ is an $R$-projective $R$-subalgebra $\Lambda$ of $A$ such that $K\Lambda=A$. We study differential graded $K$-algebras which are semisimple as $K$-algebras and define differential graded $R$-orders as a differential graded $R$-subalgebras, which are in addition classical $R$-orders in $A$. We give a series of examples for such differential graded algebras and orders. We show that any differential graded $R$-order is contained in a maximal differential graded order. We develop parts of the classical ring theory in the differential graded setting, in particular the properties of analogues of the Jacobson radical. We further define class groups of differential graded orders as subgroups of the Grothendieck group of locally free differential graded modules. We define idèles in this setting showing that these idèle groups are isomorphic to the differential graded class group. Finally we give a homomorphism to the class group of the homology of the differential graded order and prove a Mayer-Vietoris like sequence for each central idempotent of $A$, including the analogous one for the kernel groups of these morphisms.
Introduction
Differential graded algebras were introduced by Cartan in [2]. These are, by definition, a pair $(A,d)$ of an associative ${\mathbb{Z}}$-graded algebra $A$ with an endomorphism $d:A\longrightarrow A$ of degree $1$ and $d^{2}=0$, satisfying $d(a\cdot b)=a\cdot d(b)+(-1)^{|a|}a\cdot d(b)$ for all homogeneous elements $a,b\in A$. Differential graded algebras became prominent in the past few decades in order to provide a powerful tool for homological algebra.
Most research was directed in this perspective. In particular, as differential graded algebras occur naturally in the homological algebra of algebraic and differential varieties, these examples provided the guideline of reasonable assumptions. In particular, a quite systematic assumption is that differential graded algebras are connective, that is all homogeneous components in positve degrees are assumed to vanish. Often, they are algebras of infinite dimension over a base field.
We ask what happens if the algebras are more linked to classical finite dimensional algebras. Differential graded algebras are algebras at first, and so one might be tempted to understand ring theoretic properties of differential graded algebras. To our greatest surprise we could not find much research which was carried out up to now in this direction. We fist study properties such as Jacobson radicals or analogues of these, and see that the concepts in this direction in the differential graded case are different. Unlike the classical situation the intersection of all differential graded maximal left ideals is not the same as the intersection of annihilators of differential graded simple left modules. This holds however up to homotopy. Further, we prove a Nakayama’s lemma for the two-sided notion given by intersection of annihilators of simple dg-modules.
We then study differential graded orders in differential graded semisimple algebras. Since the differential vanishes automatically on central idempotents, we study finite dimensional simple algebras, i.e. matrix algebras over skew fields by Wedderburn’s theorem. We provide examples that there are various interesting differential graded structures on these matrix algebras. They bear interesting properties. In particular, their homology is not semisimple in general.
Then, we consider differential graded orders in these algebras, naively as classical orders which are stable under the differential graded structure. Graded orders were studied by LeBruyn, van Oystaen, and van den Bergh in [15]. The differential gives an important extra structure. One may ask if it would not be reasonable to ask in addition that the homology is an order in the homology of the algebra. However, this additional condition is restrictive and it does not seem to allow a rich theory.
We prove that maximal differential graded orders always exist, that any differential order is contained in a maximal one, and give examples that not all maximal orders allow a differential structure. Clearly, our algebras are not connective. Non connective differential graded algebras may have strange properties. We mention Raedschelder and Stevenson [18] for striking examples and properties in this direction.
Finally we define class groups of locally free differential graded orders. We give a $K$-theoretic definition and prove its equivalence with a second definition provided by idèles. The only modification which is necessary, is that we need to consider the subalgebra of cycles, and not the entire algebra. We then show that the functor ’taking homology’ then maps to the ordinary class group of the homology algebra, at least if the homology algebra of the ambiant semisimple algebra is semisimple. The kernel of the resulting map then bounds the fibre of stable quasi-isomorphism classes instead of stable isomorphism classes of locally free modules.
We finally mention that a differential graded Jordan-Zassenhaus theorem is not available for the moment. Hence I do not known if differential graded class groups are finite.
Here is an outline of the paper. In Section 1 we review some basic facts on differential graded algebras and differential graded modules. In Section 2 we give some basic facts on graded algebras and graded modules as far as we shall need it in particular for the homology algebra of differential graded algebras. In Section 3 we provide basically two classes of algebras which are simple as algebras and differential graded. Choosing special cases provide a wealth of interesting examples for what follows. Section 4 develops classical ring theory of differential graded algebras, in particular different concepts of radicals of differential graded algebras, providing the technical tools for discussions in further sections. Section 5 then introduces our main object of study, differential graded orders and differential graded lattices. We provide a weak version and a strong version in parallel, but focus mainly on the weak version, which seems to have more and easier accessible properties. As fist property we study maximal differential graded orders in Section 6. In Section 7 we introduce class groups of differential graded orders and differential graded idèles, show the equivalence of the K-theoretic and the idèle theoretic setting. In Section 8 we prove a Mayer-Vietoris like sequence for class groups of differential graded orders.
1. Foundations of dg-algebras and dg-modules
1.1. Generalities
Let $R$ be a commutative ring. Recall from Cartan [2] and Keller [13] that
  • a differential graded $R$-algebra (or dg-algebra for short) is a ${\mathbb{Z}}$-graded $R$-algebra $A=\bigoplus_{n\in{\mathbb{Z}}}A_{n}$ together with a graded $R$-linear endomorphism $d$ of square 00 of degree $1$ (i.e. $d(A_{n})\subseteq A_{n+1}$ and $d\circ d=0$) and such that
    $d(ab)=d(a)b+(-1)^{|a|}ad(b)$
    for all homogeneous elements $a,b\in A$. A homomorphism of differential graded algebras $f:(A,d_{A})\rightarrow(B,d_{B})$ is a degree 00 homogeneous $R$-algebra map $f$ such that $f\circ d_{A}=d_{B}\circ f$.
  • If $(A,d)$ is a differential graded algebra, then $(A^{op},d^{op})$ is a differential graded algebra (cf e.g. [21, Definition 11.1]) with $x\cdot_{op}y:=(-1)^{|x|\cdot|y|}yx$ for any homogeneous elements $x,y\in A$, and $d^{op}(x)=d(x)$. We hence write $d^{op}=d$.
  • A differential graded right $A$-module (or dg-module for short) is then a ${\mathbb{Z}}$-graded $R$-module $M$ with graded $R$-linear endomorphism $d_{M}$ of square 00 and of degree $1$, such that
    $d_{M}(ma)=d_{M}(m)a+(-1)^{|m|}md(a)$
    for all homogeneous elements $a\in A$ and $m\in M$. A differential graded $(A,d)$-left module is a differential graded $(A^{op},d)$-right module.
  • Let $(A,d_{A})$ be a differential graded $R$-algebra and let $(M,\delta_{M})$ and $(N,\delta_{N})$ be differential graded $(A,d_{A})$-modules. Then a homomorphism of differential graded modules is an $R$-linear map $f:M\rightarrow N$ with $f\circ\delta_{M}=\delta_{N}\circ f$, with $f(am)=af(m)$ for all $a\in A$ and $m\in M$.
  • If $(M,d_{M})$ is a dg-$A$-module, then let $M[k]$ be the dg-module given by $(M[k])_{n}=M_{k+n}$ for all $n\in{\mathbb{Z}}$ and $d_{M[k]}=d_{M}$. It is easy to see that $M[k]$ is a dg-module again.
  • Let $(M,d_{M})$ and $(N,d_{N})$ be differential graded $(A,d)$-modules. The homomorphism complex $Hom_{A}^{\bullet}(M,N)$ is the ${\mathbb{Z}}$-graded $A$-module given by
    $(Hom_{A}^{\bullet}(M,N))_{n}:=\{f:M\rightarrow N\;|$ $\;f\in Hom_{\mathbb{Z}}(M,N)\;\mbox{ and }$
    $f(M_{k})\subseteq N_{k{-}n}\textup{ and }$
    $f(am)=(-1)^{|a|n}a\cdot f(m)\}.$
    The elements $f$ of $Hom_{A}^{\bullet}(M,N)$ are not asked to be compatible with the differentials in any way. Let $d_{Hom}:Hom_{A}^{\bullet}(M,N)\rightarrow Hom_{A}^{\bullet}(M,N)$ given by
    $d_{Hom}(f):=d_{N}\circ f-(-1)^{|f|}f\circ d_{M}.$
    Then
    $d_{Hom}^{2}(f)$ $=$ $d_{Hom}(d_{N}\circ f-(-1)^{|f|}f\circ d_{M})$
    $=$ $d_{N}\circ(d_{N}\circ f-(-1)^{|f|}f\circ d_{M})-(-1)^{|f|-1}(d_{N}\circ f-(-1)^{|f|}f\circ d_{M})\circ d_{M}$
    $=$ $(-1)^{|f|}(d_{N}\circ f\circ d_{M}-d_{N}\circ f\circ d_{M})$
    $=$ 00 0
    Therefore $(Hom_{A}^{\bullet}(M,N),d_{Hom})$ is a complex of $R$-modules.
  • The category $A-dgMod$ of dg-modules over the dg-algebra $(A,d)$ is abelian with morphisms the degree zero cycles of $(Hom_{A}^{\bullet}(M,N),d_{Hom})$ (cf [21, Lemma 22.4.2]). If $A$ is Noetherian, then the subcategory $A-dgmod$ of finitely generated dg $A$ modules is abelian as well.
  • The homotopy category $K(A-dgMod)$ is the category with objects $A$ dg-modules and morphisms $Hom_{K(A-dgMod)}(M,N)=H_{0}(Hom_{A}^{\bullet}(M,N),d_{Hom})$.
  • If $A$ is artinian and $(A,d)$ a differential graded algebra. Then the category $A-dgmod$ has split idempotents and is hence Krull-Schmidt. This follows from the fact that the image of a homomorphism of dg-modules is a dg-module again.
Lemma 1.1.
Let $(A,d)$ be a differential graded $R$-algebra, and let $(M,d_{M})$ and $(N,d_{N})$ be differential graded $(A,d)$-modules.
  • Then $Z_{0}(Hom_{A}^{\bullet}(M,N),d_{Hom})=\ker d_{Hom}$ consists the homomorphisms of complexes, and $H_{0}(Hom_{A}^{\bullet}(M,N),d_{Hom})$ is the morphisms $M\rightarrow N$ in the homotopy category.
  • If $M=N$, then $End^{\bullet}_{A}(M)$ is a dg-algebra and composition of graded morphisms yields a differential graded $End^{\bullet}_{A}(M)$-module structure on $Hom^{\bullet}_{A}(M,N)$.
Proof: [13]
Lemma 1.2.
$d_{A}(1)=0$.
Proof. $d_{A}(1)=1\cdot d_{A}(1)+d_{A}(1)\cdot 1$, which implies $d_{A}(1)=0$.
Corollary 1.3.
Let $(A,d)$ be a differential graded $K$-algebra. Then $\ker(d)$ is a graded subalgebra of $A$.
Proof. Since $d$ is $K$-linear, $\ker(d)$ is a $K$-vector space. Since $d$ is homogeneous, $\ker(d)$ is graded. By Lemma 1.2 $\ker(d)$ contains $1$. Further, $d(xy)=d(x)y\pm xd(y)$ for any homogeneous elements $x,y$ shows that $x,y\in\ker(d)\Rightarrow xy\in\ker(d).$
Remark 1.4.
We note that Example 3.5 shows that in general $d(u)\neq 0$ for a unit $u$ in $A$. Further, units are not necessarily in degree 00, as is shown by Example 3.5.
Lemma 1.5.
Let $(A,d)$ be a differential graded algebra. Then, the algebra structure of $A$ induces a graded algebra structure on $H(A,d)$ (and hence differential graded with differential 00).
Proof. Let $a,b\in\ker(d)$. Then $d(ab)=ad(b)\pm d(a)b=0.$ Hence $ab\in\ker(d)$. Let $a=d(x)$ and $c\in\ker(d)$ be homogeneous elements. Then
$ac=d(x)c=-d(xc)-(-1)^{|b|}xd(c)=-d(xc)\in\textup{im}(d).$
Likewise $ca\in\textup{im}(d)$. Hence, the multiplication is well-defined. The fact that the additive law is well-defined and combines to a ring structure is trivial.
Lemma 1.6.
Let $(A,d)$ be a differential graded algebra and let $(M,d_{M})$ be a differential graded $A$-module. Then the $A$-module structure on $M$ induces an $H(A,d)$-module structure on $H(M,d_{M})$.
Proof. We only need to show that the induced action is well-defined.
Let $a\in\ker(d)$ be a homogeneous element, and let $m\in\ker(\delta)$ be a homogeneous element. Then
$d_{M}(m\cdot a)=d_{M}(m)\cdot a+(-1)^{|m|}m\cdot d(a)=0$
Hence $m\cdot a\in\ker(d_{M})$ again. Moreover,
$d_{M}(n)\cdot(a+d(b))$ $=$ $d_{M}(n)a+d_{M}(n)d(b)$
$=$ $d_{M}(na)-(-1)^{|a|}nd(a)+d_{M}(nd(b))-(-1)^{|b|+1}d_{M}(n)d^{2}(b)$
$=$ $d_{M}(na)+d_{M}(nd(b))-(-1)^{|a|}nd(a)$
$=$ $d_{M}(na)+d_{M}(nd(b))$
is in $\textup{im}(d_{M})$, since $a\in\ker(d)$ and hence $nd(a)=0$. Therefore, the multiplication of $H(A)$ on $H(M)$ is well-defined.
2. Group graded rings and group graded modules
2.1. Generalities on graded semisimple algebras
We recall some facts and definitions from [16]. A group graded algebra $A=\bigoplus_{g\in G}A_{g}$ is called graded simple if all non zero $G$-graded twosided ideals equal $A$. Similarly, a $G$-graded $A$-module $M$ is called graded simple if all non zero $G$-graded submodules of $M$ equal $M$. A $G$-graded $A$-module is called semisimple if $M$ is a direct sum of $G$-graded simple submodules, and a $G$-graded algebra $A$ is called semisimple if $A$ is the direct sum of $G$-graded simple algebras.
For a field $K$ artinian semisimple $G$-graded $K$-algebras are classified.
A $G$-graded $K$-algebra $A$ is $G$-graded semisimple if $A=L_{1}\oplus\dots\oplus L_{m}$ for minimal graded ideals $L_{i}$. A $G$-graded $K$-algebra $A$ is $G$-graded simple if $L_{i}=L_{j}(\sigma_{j,i})$ for some $\sigma_{j,i}\in G$ and for all $i,j$.
Theorem 2.1.
[16, 2.10.10 Theorem] (graded version of Wedderburn’s Theorem) Let $K$ be a field and let $G$ be a group. An artinian $G$-graded $K$-algebra $A$ is $G$-graded simple if and only if there is a $G$-graded division ring $D$ and an $n$-tuple of elements $(g_{1},\dots,g_{n})\in G^{n}$ such that
$A_{h}=\left(\begin{array}[]{cccc}D_{g_{1}hg_{1}^{-1}}&D_{g_{1}hg_{2}^{-1}}&\cdots&D_{g_{1}hg_{n}^{-1}}\\ D_{g_{2}hg_{1}^{-1}}&D_{g_{2}hg_{2}^{-1}}&\cdots&D_{g_{2}hg_{n}^{-1}}\\ \vdots&\vdots&&\vdots\\ D_{g_{n}hg_{1}^{-1}}&D_{g_{n}hg_{2}^{-1}}&\cdots&D_{g_{n}hg_{n}^{-1}}\\ \end{array}\right)$
where $D_{x}$ is the degree $x$ component of $D$.
Remark 2.2.
Let $K$ be a field. A finite dimensional ${\mathbb{Z}}$-graded $K$-algebra $D$ which is in addition an integral domain, is concentrated in degree 00. Indeed, if $x\neq 0$ is homogeneous of degree $d$, then $x^{n}$ is of degree $dn$, and since $D$ is an integral domain, there are infinitely many degrees ${\mathbb{N}}\cdot d$ which are of dimension at least $1$. In particular, there is no non trivial ${\mathbb{Z}}$-grading on a finite dimensional skew-field over $K$.
Note the following important result. A group graded $K$-algebra $A=\bigoplus_{g\in G}A_{g}$ is symmetric if $A\simeq A^{*}:=Hom_{K}(A,K)$ as graded $A-A$ bimodules. The $G$-grading on $A^{*}$ is given as $(A^{*})_{g}:=\{f\in A^{*}\;|\;f(A_{h})=0\forall h\neq g^{-1}\}$.
Theorem 2.3.
[5, Theorem A] A finite dimensional graded semisimple algebra is graded symmetric.
If $G=\{1\}$, then this is a classical theorem of Eilenberg and Nakayama [8, Proposition 5].
Here we are confronted with the following basic consideration. If $A$ is ${\mathbb{Z}}$-graded, then $A_{0}$ is a subalgebra, and we get that $F^{+}(A)=\bigoplus_{n\geq 0}A_{n}$, as well as $F^{-}(A)=\bigoplus_{n\leq 0}A_{n}$ are graded subalgebras of $A$. In each case, for $m>0$ the space $F_{m}^{+}(A)=\bigoplus_{n\geq m}A_{n}$ is an ideal of $F^{+}(A)$, and for any $m<0$ the space $F_{m}^{-}(A)=\bigoplus_{n\leq m}A_{n}$ is an ideal of $F^{-}(A)$. Suppose $A_{n}=0$ for $n>0$. Then $F_{1}^{+}(A)$ is a nilpotent ideal of $F^{+}(A)$, and hence part of the radical of $A$. Therefore, if $A$ is positively (or negatively) supported and $A$ is supported in only finitely many degrees, then $A$ semisimple implies $A=A_{0}$. In particular, if $A$ is semisimple artinian, and positively (or negatively) supported, then $A=A_{0}$.
Example 2.4.
  1. (1)
    Let $A=Mat_{n\times n}(D)$ for some skew-field $D$, finite dimensional over $K$, and $n>1$ an integer. Then $A$ is graded by putting $A_{0}$ the subalgebra given by the main diagonal entries
    $\left(\begin{array}[]{ccccc}\ast&0&\dots&\dots&0\\ 0&\ast&0&&\vdots\\ \vdots&\ddots&\ast&\ddots&\vdots\\ \vdots&&\ddots&\ddots&0\\ 0&\dots&\dots&0&\ast\end{array}\right)$
    Then upper one diagonal coefficients $(a_{i,i+1})_{i}$ give the degree $1$
    $\left(\begin{array}[]{ccccc}0&\ast&0&\dots&0\\ 0&0&\ast&\ddots&\vdots\\ \vdots&\ddots&\ddots&\ddots&0\\ \vdots&&\ddots&0&\ast\\ 0&\dots&\dots&0&0\end{array}\right)$
    the upper $2$ diagonal coefficients $(a_{i,i+2})_{i}$ give the degree $2$, etc. The lower one diagonal coefficients $(a_{i+1,i})_{i}$ give the degree $-1$, the lower $2$ diagonal coefficients $(a_{i+2,i})_{i}$ give the degree $-2$, etc. Elementary matrix multiplication proves the definition for a graded algebra. Moreover, $A$ is semisimple as an algebra.
  2. (2)
    Let $A=\bigoplus_{n\in{\mathbb{Z}}}A_{n}$ be a graded semisimple algebra, then for any $u\in A^{\times}$ we see that $A^{u}:=\bigoplus_{n\in{\mathbb{Z}}}uA_{n}u^{-1}$ is again a graded semisimple algebra.
Lemma 2.5.
Let $A$ be a graded semisimple ${\mathbb{Z}}$-graded algebra. If $\text{rad}(A_{0})\cdot A_{n}=A_{n}\cdot\text{rad}(A_{0})$, and if $A_{0}$ is artinian, then $A_{0}$ is semisimple as well.
Proof. Let $A=\bigoplus_{n\in{\mathbb{Z}}}A_{n}$ and let $\widetilde{I}=\text{rad}(A_{0})$. Since $A_{0}$ is artinian, $\text{rad}(A_{0})$ is a nilpotent ideal. Moreover, $\text{rad}(A_{0})\cdot A_{n}=A_{n}\cdot\text{rad}(A_{0})$ implies that $\text{rad}(A_{0})A=\bigoplus_{n\in{\mathbb{Z}}}\text{rad}(A_{0})A_{n}$ is a graded nilpotent ideal of $A$, whence included in $\text{rad}(A)$. Since $A$ is semisimple, $\text{rad}(A)=0$ and hence $\text{rad}(A_{0})=0$. This completes the proof.
2.2. Semisimplicity in the category of differential graded algebras
We recall a theorem of Tempest and Garcia-Rozas [1]. A differential graded $A$-module $X$ is simple if the only differential graded $A$-submodules are 00 and $X$. A differential graded $A$-module $X$ is semisimple if $X$ is a direct sum of simple differential graded $A$-modules.
Theorem 2.6.
  • [1, Theorem 4.7] Let $(A,d)$ be a differential graded algebra. Then the following are equivalent:
    1. (1)
      The regular differential graded $A$-module $A$ is projective in the category of dg-modules
    2. (2)
      $A$ is acyclic.
    3. (3)
      $1_{A}\in im(d)$
    4. (4)
      any left dg-module is acyclic.
    5. (5)
      The functor $A\otimes_{Z(A)}-$ from graded $Z(A)$-modules to differential graded $A$-modules is an equivalence with quasi-inverse $Z(-)$.
  • [1, Theorem 5.3] A differential graded algebra $(A,d)$ is semisimple precisely when the graded algebra of cycles $Z(A)=\ker(d)$ is semisimple and $A$ is an acyclic complex. This in turn is equivalent with the fact that all differential graded $A$-modules are semisimple (as differential graded $A$-modules).
Note that Tempest and Garcia-Rochas show [1, Lemma 4.2] that if $1_{A}=d(z)$, then $A=Z(A)\oplus Z(A)y$.
Remark 2.7.
Let $(A,d)$ be a differential graded $K$-algebra, semisimple as algebra, and let $e^{2}=e\in Z(A)$. Then, copying the argument of [4, Theorem 1.4 and Corollary 1.5], since $e^{2}=e\in Z(A)$, we get that $e$ is necessarily in $A_{0}$. Moreover,
$d(e)=d(e^{2})=d(e)e+ed(e)=2ed(e)=4ed(e)$
Hence $ed(e)=0$. Moreover $(1-e)d(e)=(1-e)d(e)e+(1-e)ed(e)=0$, and hence $d(e)=0$. Therefore, if $A=A_{1}\times A_{2}$ as algebra, then $(A,d)=(A_{1},d)\times(A_{2},d)$ as differential graded algebras.
Corollary 2.8.
As we have seen in Remark 2.7 for a differential graded algebra $A$ which is semisimple artinian as algebra, by Wedderburn’s theorem we have that $A$ is a direct sum of differential graded algebras, each of which is a matrix algebra of a skew field.
3. Some examples
We produce a series of examples showing that our concept is non trivial and allows interesting phenomena.
Remark 3.1.
Note that Example 2.4 provides gradings such that $A\cdot e\not\subseteq A\cdot d(e)$ for any primitive idempotent $e$ of degree 00 and any differential $d$ of degree $1$. Also for $End_{K}^{\bullet}(L)$ for a bounded complex $L$ of $K$-vector spaces there is such an idempotent. Indeed, in this case the degree 00 component is formed by matrix algebras along the diagonal, and the degree $1$ component is formed by matrices right up to these.
Proposition 3.2.
Let $K$ be a field and let $(A,d)$ be a finite dimensional differential graded algebra. Suppose that $A$ is a split simple $K$-algebra. Suppose that there is a primitive idempotent $e$ of $A$ such that $A\cdot e\not\subseteq A\cdot d(e)$. Then there is a bounded complex $L$ of $K$-modules such that $A\simeq Hom_{K}^{\bullet}(L,L)^{op}$ as differential graded algebras. Conversely, $A=Hom_{K}^{\bullet}(L,L)$ is differential graded, finite dimensional simple as algebra, such that there is a primitive idempotent $e$ with $A\cdot e\not\subseteq A\cdot d(e)$.
Proof. If $L$ is a bounded complex of $K$-vector spaces, then $Hom_{K}^{\bullet}(L,L)$ is a full matrix ring over $K$, as ungraded algebra, and hence simple as algebra. Further, $Hom_{K}^{\bullet}(L,L)$ is a differential graded algebra by Lemma 1.1.
Conversely, let $K$ be a field and let $(A,d)$ be a finite dimensional differential graded algebra. Suppose that $A$ is a split simple $K$-algebra. By Wedderburn’s theorem, $A$ is a full matrix algebra over $K$. Let $e$ be a primitive idempotent of $A$ satisfying $Ae\not\subseteq Ad(e)$. Then
$M:=A\cdot e+A\cdot d(e)\textup{ and }N:=A\cdot d(e)$
are differential graded $(A,d)$-modules. Further $N<M$ and
$L:=M/N\neq 0$
is a differential graded $(A,d)$-module. As $A$-module, we see that $L\simeq Ae$ is a progenerator. Hence $L$ is a natural differential graded $(A,d)-(End^{\bullet}(L),d_{Hom})$ bimodule. Now, for any homogeneous $a\in A$, left multiplication by $a$ gives a homogeneous element $\varphi(a)\in End^{\bullet}(L)$. Further, $\varphi$ is additive, sends $1\in A$ to the identity on $L$, and induces a ring homomorphism
$\varphi:A\longrightarrow End^{\bullet}(L)^{op}.$
Since $L$ is a progenerator, $\varphi$ is injective. Since $\dim_{K}(A)=\dim_{K}(End^{\bullet}(L))$, we get that $\varphi$ is an isomorphism of algebras. Now, for any homogeneous $a,b\in A$, we have
$d(a)b=d(ab)-(-1)^{|a|}ad(b)$
we get
$\varphi(d(a))=d\circ\varphi(a)-(-1)^{|a|}\varphi(a)\circ d=d_{Hom}(\varphi(a))$
and therefore $\varphi$ is an isomorphism of differential graded algebras. Further, by Remark 3.1 there is a primitive idempotent $e$ of degree 00, by degree considerations we get $Ae\not\subseteq Ad(e)$.
Remark 3.3.
The hypothesis that there is a primitive idempotent $e$ of $A$ such that $A\cdot e\not\subseteq A\cdot d(e)$ is superfluous. Indeed, [4, Corollary 1.5] shows that $A$ is isomorphic, as a graded algebra, to a matrix algebra with the main diagonal in degree 00. Transporting the differential structure via this isomorphism, we obtain the existence of such an idempotent.
Remark 3.4.
After having finished and submitted the manuscript I discovered that D. Orlov defined our notion of simple differential graded algebra earlier in [17], and called it abstractly simple. Moreover, he proved the statement of Proposition 3.2 by completely different means. His proof uses scheme theoretic arguments. However, he has to assume that the primitive central idempotents are in degree 00. Our approach however gives that this can be assumed to be automatically satisfied using [4, Corollary 1.5].
Example 3.5.
We consider a field $K$ and the algebra of $2\times 2$ matrices over $K$. We shall use Proposition 3.2. Except the stalk complex and complexes with differential 00, the only possible complex realising this dg-algebra is
$\cdots\longrightarrow 0\longrightarrow K\stackrel{{\scriptstyle\cdot x}}{{\longrightarrow}}K\longrightarrow 0\longrightarrow\cdots.$
Its endomorphism algebra is $Mat_{2}(K)$ with a grading as in Example 2.4.1. The differential is given by
$d(\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right))=\left(\begin{array}[]{cc}x&0\\ 0&x\end{array}\right)$
and
$d(\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right))=d(\left(\begin{array}[]{cc}0&0\\ 0&-1\end{array}\right))=\left(\begin{array}[]{cc}0&-x\\ 0&0\end{array}\right).$
This gives a dg-algebra and any differential is of this form for some $x\in K$. If $x\neq 0$, the homology is 00.
Note that if $x\neq 0$, then the kernel of the differential is
$Z(A)=\{\left(\begin{array}[]{cc}a&b\\ 0&a\end{array}\right)\;|\;a,b\in K\}\simeq K[\epsilon]/\epsilon^{2}$
with $\epsilon$ in degree $1$. This is not semisimple, and hence we do not get a semisimple differential graded algebra in the sense of Aldrich and Garcia-Rochas [1]. However, for $x$ invertible we have $z=\frac{1}{x}\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right)$ is a preimage of $1$, as required in Aldrich and Garcia-Rochas [1] structure theorem.
We further observe that $\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)$ is invertible, non homogeneous with a summand of degree $-1$ and another summand of degree $1$. However,
$d(\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right))=d(\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right))+d(\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right))=\left(\begin{array}[]{cc}x&0\\ 0&x\end{array}\right)$
is non zero for $x\neq 0$. Hence, the differential of invertible elements are not necessarily 00. Nevertheless, $d(1)=0$.
As any differential graded $(A,d)$-module $(M,\delta)$ is at first an $A$-module, it is useful to consider the possible dg-module structures on $K^{2}$ for the above dg-algebra $(Mat_{2}(K),d_{x})$. So, let $\delta(\left(\begin{array}[]{c}0\\ 1\end{array}\right))=\left(\begin{array}[]{c}u\\ v\end{array}\right)$ for some $u,v\in K$ and $\delta(\left(\begin{array}[]{c}1\\ 0\end{array}\right))=\left(\begin{array}[]{c}s\\ t\end{array}\right)$ for some $s,t\in K$.
00 0 $=$ $\delta(\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right)\cdot\left(\begin{array}[]{c}0\\ 1\end{array}\right))=\left(\begin{array}[]{cc}x&0\\ 0&x\end{array}\right)\cdot\left(\begin{array}[]{c}0\\ 1\end{array}\right)-\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right)\cdot\left(\begin{array}[]{c}u\\ v\end{array}\right)=\left(\begin{array}[]{c}0\\ x-u\end{array}\right)$
00 0 $=$ $\delta(\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right)\cdot\left(\begin{array}[]{c}0\\ 1\end{array}\right))=\left(\begin{array}[]{cc}0&-x\\ 0&0\end{array}\right)\cdot\left(\begin{array}[]{c}0\\ 1\end{array}\right)+\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right)\cdot\left(\begin{array}[]{c}u\\ v\end{array}\right)=\left(\begin{array}[]{c}u-x\\ 0\end{array}\right)$
$\left(\begin{array}[]{c}u\\ v\end{array}\right)$ $=$ $\delta(\left(\begin{array}[]{cc}0&0\\ 0&1\end{array}\right)\cdot\left(\begin{array}[]{c}0\\ 1\end{array}\right))=\left(\begin{array}[]{cc}0&x\\ 0&0\end{array}\right)\cdot\left(\begin{array}[]{c}0\\ 1\end{array}\right)+\left(\begin{array}[]{cc}0&0\\ 0&1\end{array}\right)\cdot\left(\begin{array}[]{c}u\\ v\end{array}\right)=\left(\begin{array}[]{c}x\\ v\end{array}\right)$
$\left(\begin{array}[]{c}s\\ t\end{array}\right)$ $=$ $\delta(\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)\cdot\left(\begin{array}[]{c}0\\ 1\end{array}\right))=-\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)\cdot\left(\begin{array}[]{c}u\\ v\end{array}\right)=\left(\begin{array}[]{c}v\\ 0\end{array}\right)$
$\left(\begin{array}[]{c}u\\ v\end{array}\right)$ $=$ $\delta(\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right)\cdot\left(\begin{array}[]{c}1\\ 0\end{array}\right))=\left(\begin{array}[]{cc}x&0\\ 0&x\end{array}\right)\cdot\left(\begin{array}[]{c}1\\ 0\end{array}\right)-\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right)\cdot\left(\begin{array}[]{c}s\\ t\end{array}\right)=\left(\begin{array}[]{c}x\\ s\end{array}\right)$
$\left(\begin{array}[]{c}as\\ 0\end{array}\right)$ $=$ $\delta(\left(\begin{array}[]{cc}a&0\\ 0&b\end{array}\right)\cdot\left(\begin{array}[]{c}1\\ 0\end{array}\right))=\left(\begin{array}[]{cc}0&x(b-a)\\ 0&0\end{array}\right)\cdot\left(\begin{array}[]{c}1\\ 0\end{array}\right)+\left(\begin{array}[]{cc}a&0\\ 0&b\end{array}\right)\cdot\left(\begin{array}[]{c}s\\ t\end{array}\right)=\left(\begin{array}[]{c}as\\ bt\end{array}\right)$
00 0 $=$ $\delta(\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)\cdot\left(\begin{array}[]{c}1\\ 0\end{array}\right))=-\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)\cdot\left(\begin{array}[]{c}s\\ t\end{array}\right)=\left(\begin{array}[]{c}-t\\ 0\end{array}\right)$
00 0 $=$ $\delta^{2}(\left(\begin{array}[]{c}0\\ 1\end{array}\right))=\delta(\left(\begin{array}[]{c}x\\ v\end{array}\right))=\left(\begin{array}[]{c}vx\\ v^{2}\end{array}\right)+\left(\begin{array}[]{c}vx\\ 0\end{array}\right)=\left(\begin{array}[]{c}2vx\\ v^{2}\end{array}\right)$
Hence $s=v=t=0$, $x=u$. Since $\delta$ is of degree $1$, we need to have $\left(\begin{array}[]{c}1\\ 0\end{array}\right)$ is of degree $n$ and $\left(\begin{array}[]{c}0\\ 1\end{array}\right)$ is of degree $n-1$ for some integer $n$. Up to shift there is a unique dg-module structure given by $\delta(\left(\begin{array}[]{c}p\\ q\end{array}\right))=\left(\begin{array}[]{c}xq\\ 0\end{array}\right)$. Of course, this is a simple differential graded module.
Let $(C^{*},\widehat{\delta})$ be a bounded complex of finitely generated $Mat_{2\times 2}(K)$-modules. Each homogeneous component $C^{n}$ is isomorphic to a a direct sum of differential graded modules $(K^{2},\delta)$ as above. Then, the total complex is a differential graded module. Using the above explicit computation can be used to show that this gives again the only possible differential graded structures on the corresponding graded module.
Remark 3.6.
I wish to thank Bernhard Keller who noted111email from December 24, 2022 to the author that the above differential graded algebra is the differential graded endomorphism algebra $Hom_{K}^{\bullet}(M,M)$ for $M$ being the complex
$\cdots\longrightarrow 0\longrightarrow K\stackrel{{\scriptstyle\cdot x}}{{\longrightarrow}}K\longrightarrow 0\longrightarrow\cdots.$
Example 3.7.
We consider the $3\times 3$ matrix ring over a commutative ring $R$. We will use an alternative grading coming from the fact that this ring is Morita equivalent to the $2$ matrix ring over $R$.
$Mat_{3}(R)=\{\left(\begin{array}[]{ccc}a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ a_{31}&a_{32}&a_{33}\end{array}\right)\;$ ||| $\;{deg(a_{31})=deg(a_{32})=}-1;$
$deg(a_{11})=deg(a_{12})=deg(a_{21})=deg(a_{22})=deg(a_{33})=0;$
$deg(a_{13})=deg(a_{23})=1\}$
This grading comes from the complex $L$
$\cdots\longrightarrow 0\longrightarrow R\stackrel{{\scriptstyle a_{11}\choose a_{21}}}{{\longrightarrow}}R^{2}\longrightarrow 0\longrightarrow\cdots$
We make explicit the differentials.
We get
$d(\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ x&y&0\end{array}\right))=\left(\begin{array}[]{ccc}a_{11}x&a_{11}y&0\\ a_{21}x&a_{21}y&0\\ 0&0&a_{11}x+a_{21}y\end{array}\right)$
and
$d(\left(\begin{array}[]{ccc}\lambda&\sigma&0\\ \tau&\mu&0\\ 0&0&\nu\end{array}\right))=\left(\begin{array}[]{ccc}0&0&a_{11}(\nu-\lambda)-a_{21}\sigma\\ 0&0&a_{21}(\nu-\mu)-a_{11}\tau\\ 0&0&0\end{array}\right)$
We may study the homology of this differential graded algebra. Suppose that $R$ is an integral domain. If $a_{11}\neq 0\neq a_{21}$, then the degree 00 cycles is given by the set of matrices
$\left(\begin{array}[]{ccc}\lambda&\frac{a_{11}}{a_{21}}(\nu-\lambda)&0\\ \frac{a_{21}}{a_{11}}(\nu-\mu)&\mu&0\\ 0&0&\nu\end{array}\right)$
for $\lambda,\mu,\nu\in R$. This forms a $3$-dimensional affine variety. The boundaries in degree 00 forms a $2$-dimensional subvariety. Hence, the degree 00 homology has an $R$-torsion free part of rank $1$. Degree $-1$ cycles are trivial, since $R$ has no 00 divisors. The degree $1$ cycles is $R^{2}$, by definition. Clearly, the degree $1$ homology is $\left(R/(a_{11}R+a_{21}R)\right)^{2}$.
4. Dg-algebras with respect to Semisimplicity and Noetherian ring theoretical properties
As we shall need to consider semisimplicity of differential graded algebras, it makes sense to consider differential graded versions of the Jacobson radicals. Furthre, we shall consider semisimplicity of the homology algebra.
4.1. Semisimplicity of the homology
Example 4.1.
Consider Example 3.7. The special case $a_{11}=0\neq a_{21}$ produces a dg-algebra $A$, which is semisimple as algebra, but $H(A)=K^{2}\oplus K[1]$ as graded vector space, i.e. $H(A)$ is a $3$-dimensional algebra. $H(A)$ is $2$-dimensional in degree 00, actually isomorphic to the algebra of dual numbers over $K$, and $1$-dimensional in degree $-1$. More precisely, as algebra, $H(A)\simeq K[X,Y]/(XY-YX,X^{2},Y^{2})$, where $X$ is in degree $-1$ and $Y$ is in degree 00. In particular, $H(A)$ is not semisimple as algebra.
In this context is may be worth to mention the following criterion in this context.
Lemma 4.2.
Let $K$ be a field and let $(A,d)$ be a finite dimensional differential graded $K$-algebra. Then $(\ker(d),d)$ is a subalgebra of $(A,d)$. If $\ker(d)$ is semisimple, then $H(A,d)$ is semisimple as well. Moreover, the restriction of the epimorphism $\ker(d)\longrightarrow\kern-8.00003pt\rightarrow H(A,d)$ to $\ker(d)^{\times}$ yields a surjective group homomorphism $\ker(d)^{\times}\longrightarrow H(A,d)^{\times}$.
Proof. By Corollary 1.3 $\ker(d)$ is a graded subalgebra of $A$. By hypothesis, $\ker(d)$ is a semisimple $K$-algebra. Now, by Lemma 1.5 there is an epimorphism $\ker(d)\stackrel{{\scriptstyle\pi}}{{\longrightarrow\kern-8.00003pt\rightarrow}}H(A,d)$ of algebras, and the preimage of a twosided ideal in $H(A,d)$ is a twosided ideal of $\ker(d)$. Therefore $H(A,d)\times\ker(\pi)=\ker(d)$ as algebras, and as a consequence any unit $H(A,d)$ lifts to a unit of $\ker(d)$.
Remark 4.3.
Let $(A,d)$ be a semisimple artinian dg-algebra and let $(L,d_{L})$ and $(M,d_{M})$ be differential graded $(A,d)$-modules. Then $\ker({(d_{Hom}})_{0})=End_{(A,d)}(L,d)$ is not semisimple in general since not every homomorphism $\alpha:(L,d_{L})\rightarrow(M,d_{M})$ of complexes decomposes as $\alpha=\gamma\circ\delta$ for a split epimorphism $\delta$ of complexes and a split monomorphism $\gamma$ of complexes.
We recall the easy argument. We consider the special case when $A$ is concentrated in degree 00. Since $A$ is semisimple, $L$ is a semisimple $A$-module. We claim that $L=L_{0}\oplus H(L)$ is the direct sum of contractible $2$-term complexes $L_{0}$ and its homology $H(L)$. Indeed, since there is no non zero $A$-module homomorphism between two non isomorphic simple $A$-modules, we may assume, multiplying by a specific central idempotent if necessary, that $A$ actually is simple. By Morita equivalence we can therefore assume that $A$ is a skew field. Then, we may choose a basis $B_{0}$ of $L_{0}$ and then choosing a well adapted basis $B_{1}$ in degree $1$ and a well adapted basis in degree $-1$, we may assume that $d_{-1}(b)\in B_{0}\cup\{0\}$ for any $b\in B_{-1}$, and using that $d^{2}=0$, likewise $d_{0}(b)\in B_{1}\cup\{0\}$ for any $b\in B_{0}$. Iteratively, modifying the already chosen bases if necessary on those parts which are in the kernel/cokernel of the differentials one degree above/below, we may assume that $d(b)\in B_{n}\cup\{0\}$ for any $b\in B_{n-1}$ for all $n$.
Consider the case of
[…]
concentrated in degree 00 and $1$, and
[…]
concentrated in degree $-1$ and 00. Moreover, suppose that $A=D$ is a skew field. Then $Hom_{A}(L_{1},L_{2})=Hom_{D}(M,N)$ from the spaces in degree 00, using that the differentials do not impose any restriction. However, $Hom_{A}((L_{2},L_{1})=0$ since the compatibility of the differentials impose that the map in degree 00 needs to be 00.
Remark 4.4.
Consider the situation of Remark 4.3. Then we have a ring homomorphism
${End_{A}^{\bullet}(L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{0}(End_{A}^{\bullet}(L))=End_{gr-A}(H(L))}$
from the endomorphism complex of $L$ to the ring of graded $A$-linear homomorphisms of $H(L)$. Since $A$ is semisimple, and since $H(L)$ is a semisimple $A$-module, $End_{gr-A}(H(L))$ is a semisimple algebra. Moreover, using Lemma 1.1, we get that $(End_{A}^{\bullet}(L),d_{Hom})$ is a differential graded algebra, mapping onto a semisimple algebra $End_{A}(H(L))$.
4.2. Differential graded radicals of differential graded algebras
We shall study the ring theory of differential graded algebras.
Lemma 4.5.
Let $(\Lambda,d)$ be a differential graded algebra, let $(M,\delta)$ be a differential graded $(\Lambda,d)$-module with a differential graded submodule $(N,\delta)$. Then there is a differential $\overline{\delta}$ on $M/N$ given by $\overline{\delta}(m+N):=\delta(m)$ for all $m\in M$.
Proof. $M/N$ is a $\Lambda$-module by general ring theory. Since $\delta(N)\subseteq N$, the map $\overline{\delta}$ is well-defined. $\overline{\delta}$ is clearly additive since $\delta$ is additive. Since $\delta^{2}=0$, also $\overline{\delta}^{2}=0$. Since $\delta$ satisfies Leibniz formula, so is $\overline{\delta}$.
Lemma 4.6.
Let $(\Lambda,d)$ be a differential graded algebra and let $(I,d)$ be a differential graded ideal of $(\Lambda,d)$. Then there is a maximal differential graded ideal $(M,d)$ in $(\Lambda,d)$ containing $(I,d)$.
Proof. An analogue of the usual proof in the classical situation works. We use Zorn’s lemma. Let $\mathcal{X}$ be the set of differential graded ideals of $(\Lambda,d)$ containing $(I,d)$ but different from $(\Lambda,d)$. Then $(I,d)\in\mathcal{X}$, and hence $\mathcal{X}$ is not empty. Let ${\mathcal{L}}$ be a totally ordered subset of $\mathcal{X}$ and let $J:=\bigcup_{L\in{\mathcal{L}}}L$. Then $J$ is an ideal by the classical argument. Moreover, it is differential graded since if $x\in J$, there is $L_{x}\in\mathcal{L}$ with $x\in L_{x}$. Then $d(x)\in L_{x}\subseteq J$. Further the Leibniz formula holds in $\Lambda$, whence also in $J$. If $1\in J$, then there is $L_{1}\in\mathcal{L}$ such that $1\in L_{1}$, and hence $L_{1}$ is not in $\mathcal{X}$, which provides a contradiction since ${\mathcal{L}}\subseteq{\mathcal{X}}$.
Lemma 4.7.
Let $(\Lambda,d)$ be a differential graded algebra and let $(I,d)$ be a differential graded ideal of $(\Lambda,d)$. Then $(I,d)$ is a maximal differential graded ideal if and only if $(\Lambda/I,\overline{d})$ does not contain non trivial differential graded ideals, for $\overline{d}$ the differential induced on $\Lambda/I$ by $d$.
Proof. Let $x\in\Lambda$ and $y\in I$. Then by Lemma 4.5 for $\overline{d}(x+I):=d(x)+I$ defines a differential on $\Lambda/I$. Consider the zero ideal in $(\Lambda/I,\overline{d})$. By Lemma 4.6 this is contained in a maximal differential graded ideal $(\overline{M},\overline{d})$. Let $M$ be the preimage of $\overline{M}$ in $\Lambda$. We claim that $(M,d)$ is a dg-ideal. Since $\overline{M}$ is an ideal, by classical ring theory $M$ is an ideal as well. For $m\in M$ let $d(m)+I=\overline{d}(m+I)\in\overline{M}$ since $(\overline{M},\overline{d})$ is a dg-ideal. Hence $d(m)\in M$ and by consequence $(M,d)$ is a differential graded ideal. If there is no non trivial differential graded ideal in $\Lambda/I$, then $M=I$ and $I$ is a maximal differential graded ideal. If $\Lambda/I$ admits non trivial differential graded ideals, then $I\neq M$, and $I$ is not maximal. This shows the statement.
Remark 4.8.
Note that the statement of Lemma 4.7 is completely formal and holds for left- or right- or twosided ideals.
Lemma 4.9.
Let $(\Lambda,d)$ be a differential graded algebra. Then for any homogeneous $a\in ker(d)$ we have that $\Lambda\cdot a$ is a differential graded $(\Lambda,d)$ left ideal and $a\cdot\Lambda$ is a differential graded $(\Lambda,d)$ right ideal.
Proof. Indeed, $d(\lambda\cdot a)=d(\lambda)\cdot a\pm\lambda d(a)=d(\lambda)\cdot a$ and likewise $d(a\cdot\lambda)=d(a)\lambda\pm a\cdot d(\lambda)=a\cdot d(\lambda)$.
Remark 4.10.
If $a$ is not (necessarily) a cycle, i.e. $a\not\in ker(d)$, then $\Lambda\cdot a+\Lambda\cdot d(a)$ is a differential graded ideal. Similarly, if $(M,d_{M})$ is a dg-bimodule, and if $m\in\ker(d_{M})$, then $\Lambda\cdot m+\Lambda\cdot d_{M}(m)$ is a left dg-submodule of $(M,d_{M})$.
Corollary 4.11.
Let $(\Lambda,d)$ be a differential graded algebra and let $(I,d)$ be a maximal left (resp. right) differential graded ideal. If $a\in(ker(d)+I)$ is homogeneous, and if $a\not\in I$, then $a+I$ has a left (resp. right) inverse in $\Lambda/I$.
Proof. If $a\in ker(d)+I$, but $a\not\in I$, then $a+I\in ker(\overline{d})$ and then there is $x\in I$ such that $a+x\in\ker(d)$ and hence $\Lambda\cdot(a+x)$ is a differential graded ideal of $\Lambda$. Hence, also $I+\Lambda(a+x)=I+\Lambda a$ is a differential graded ideal of $\Lambda$. If $a\not\in I$, then by the maximality of $(I,d)$ we have that $\Lambda=I+\Lambda a$. Therefore $a+I$ has a left inverse in $\Lambda/I$. The statement for the right ideals is analogous.
Definition 4.12.
Let $(\Lambda,d)$ be a differential graded algebra. A non zero differential graded $(\Lambda,d)$-module $(M,\delta)$ is called differential graded simple if there is no differential graded submodule of $(M,\delta)$ different from $(M,\delta)$ and 00. A differential graded $(\Lambda,d)$-module $(M,\delta)$ is called differential graded semisimple if $(M,\delta)$ is the direct sum of simple differential graded $(\Lambda,d)$-modules. A differential graded algebra $(\Lambda,d)$ is called semisimple if $(\Lambda,d)$ as left $(\Lambda,d)$-module is semisimple.
Example 4.13.
Let $K$ be a field and let $A=K[X]/X^{2}$ for $deg(X)=-1$. Then $d(1)=0$ and $d(X)=1$ give a structure of differential graded algebra to $A$. Indeed, $(a+bX)(c+dX)=ab+(bc+ad)X$ and
$d((a+bX)(c+dX))=bc+ad$
and
$d(a+bX)\cdot(c+dX)+(a-bX)\cdot d(c+dX)=b(c+dX)+(a-bX)d=bc+ad.$
Further, $00=d(0)=d(X^{2})=d(X)\cdot X-X\cdot d(X)=X-X=$ 0. The algebra $A$ is simple as differential graded algebra. Indeed, there is only one non trivial algebra ideal, namely $XK[X]$. However, this is not a differential graded ideal, since $d(X)=1$. Further, the only simple differential graded module is the regular one, which is not concentrated in a single degree. Note that this is the smallest example of algebras mentioned in the second item of Theorem 2.6 studied by Aldrich and Garcia Rozas [1]. Of course, $A$ is not a simple algebra.
Corollary 4.14.
Let $(\Lambda,d)$ be a differential graded algebra and let $(I,d)$ be maximal differential graded left ideal of $(\Lambda,d)$. Then $(\Lambda/I,\overline{d})$ is a simple differential graded $(\Lambda,d)$-module.
Proof. This is a direct consequence of Lemma 4.7.
Lemma 4.15.
Let $(\Lambda,d)$ be a differential graded algebra and let $(S,\delta)$ be a simple differential graded $(\Lambda,d)$-module. Then for any non zero homogeneous $u\in\ker(\delta)$ there is a surjective homomorphism $(\Lambda,d)[|u|]\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}(S,\delta)$ given by $\pi(\lambda)=\lambda u$. Further, there is always a non zero homogeneous $u\in\ker(\delta)$, and $\ker(\pi)$ is a maximal differential graded ideal.
Proof. Due to the shift, $\pi$ is a homomorphism of graded modules. We need to show that $\pi$ is compatible with the differential. Let $\lambda\in\Lambda$ be homogeneous. Then
$\pi(d(\lambda))=d(\lambda)u=d(\lambda)u+(-1)^{|\lambda|}\lambda\delta(u)=\delta(\lambda u)=\delta(\pi(\lambda)),$
where the third equation is Leibniz formula, and where the second equation holds since $u\in\ker\delta$. Using Lemma 4.9 we see that $\textup{im}(\pi)$ is a non trivial dg submodule of $S$, and by simplicity of $S$, we see that $\pi$ is surjective. Note that $\delta^{2}=0$ shows that $\ker(\delta)$ is not zero. Using Lemma 4.7 we see that $\ker(\pi)$ is a maximal dg-ideal.
Lemma 4.16.
Let $(\Lambda,d)$ be a differential graded algebra, let $(M,\delta)$ be a differential graded $(\Lambda,d)$-module. Suppose that there are simple differential graded $(\Lambda,d)$-left modules $(S_{1},\delta_{1})$, …, $(S_{n},\delta_{n})$ such that
$(M,\delta)\simeq(S_{1},\delta_{1})\oplus\dots\oplus(S_{n},\delta_{n})$
as differential graded left modules. Then for any differential graded submodule $(N,\delta)$ of $(M,\delta)$ there is a differential graded submodule $(L,\delta)$ of $(M,\delta)$ such that $M=N\oplus L$.
Proof. Consider the family of sets $\{i_{1},\dots,i_{k}\}\subseteq\{1,\dots,n\}$ such that $\bigoplus_{j=1}^{k}S_{i_{j}}\cap N=\{0\}$. Then taking a maximal one $P=\{i_{1},\dots,i_{k}\}$ of these subsets, and put $L:=\bigoplus_{j=1}^{k}S_{i_{j}}$. Then $L+N=L\oplus N$ is a differential graded $(\Lambda,d)$-module. Hence we get that $(S_{\ell}+L)\cap N\neq\{0\}$ for any $\ell\not\in P$. Therefore $(L\oplus N)\cap S_{\ell}\neq\{0\}$ for any $\ell\not\in P$. Since $S_{\ell}$ is a differential graded simple $(\Lambda,d)$-module, and since $L\oplus N$ is a differential graded module, $S_{\ell}\subseteq L\oplus N$ for any $\ell\not\in P$. But this shows that $L\oplus N\supseteq\bigoplus_{i=1}^{n}S_{i}=M$. We proved the statement.
Definition 4.17.
Let $K$ be a field and let $(A,d)$ be a differential graded $K$-algebra. The dg-algebra $A$ is called differential graded simple if the only differential graded twosided $(A,d)$-ideals are 00 and $(A,d)$.
Remark 4.18.
Recall from the second item of Theorem 2.6 that every differential graded $(A,d)$-module is semisimple if and only if $(A,d)$ is exact and $\ker(d)$ is a semisimple algebra. Moreover, then $A=\ker(d)\oplus z\cdot\ker(d)$ for $z\in d^{-1}(1)$.
Consider for example the case of an algebra concentrated in degree 00, such as the field $K$. Then there is no non trivial twosided ideal. However, this dg-algebra does not satisfy the condition of Theorem 2.6. The classical result that simple artinian algebras are semisimple cannot be transposed to the dg-situation.
Proposition 4.19.
Let $K$ be a field and let $(A,d)$ be a differential graded algebra without any differential graded twosided ideal other than 00 and $A$. Suppose that the regular differential graded left $(A,d)$ module and the regular differential graded right $(A,d)$ module are artinian.
Denote by $\textup{cone}(id_{K})$ the cone of the identity map of the stalk complex $K$ in $D^{b}(K-mod)$.
Then, as differential graded left $(A,d)$-module, the regular module $(A,d)$ is the direct sum of shifts of copies of one specific simple differential graded modules $(S,\delta)$ and shifted copies of modules of the form total complex of the tensor product of $(S,\delta)$ by $\textup{cone}(id_{K})$, i.e. $tot\left((S,\delta)\otimes_{K}(0\longrightarrow K\stackrel{{\scriptstyle id}}{{\longrightarrow}}K\longrightarrow 0)\right)$.
Proof. As $A$ is artinian there is a minimal non zero differential graded left ideal $(S,d)$ and a minimal non zero differential graded right ideal $(T,d)$. Then consider the $(A,d)-(A,d)$-bimodule $X:=(S,d)\otimes_{K}(T,d)$. The multiplication map
$A\otimes_{K}A\longrightarrow A$
is a chain map (replacing $A\otimes_{K}A$ by $tot(A\otimes_{K}A)$) by the Leibniz rule and maps $X$ to a bimodule, whence a twosided ideal of $A$. Further, it is generated by $S\cdot a$ for some $a\in\ker(d)$ and $S\cdot a+S\cdot d(a)$ for some $a\not\in\ker(d)$. Hence the image is a differential graded twosided ideal of $(A,d)$. Since there is no non trivial twosided ideal, this image is $A$. Now, consider $X$ as differential graded left module. Since only the $K$-structure of $(T,d)$ is of importance then, and since complexes of $K$-modules are a direct sum of shifted copies of stalk complexes $K$ and shifted copies of complexes $\textup{cone}(id_{K})$. This shows the statement.
Remark 4.20.
We should note that the short exact sequence
[…]
of differential graded $K$-modules yields a short exact sequence
$0\longrightarrow(S,\delta)\longrightarrow tot\left((S,\delta)\otimes_{K}(0\longrightarrow K\stackrel{{\scriptstyle id}}{{\longrightarrow}}K\longrightarrow 0)\right)\longrightarrow(S,\delta)[1]\longrightarrow 0$
of $(A,d)$-left modules.
For any differential graded left $(\Lambda,d)$-module $(M,\delta)$ denote the annihilator of $(M,\delta)$ by
$\textup{ann}(M,\delta):=\{\lambda\in\Lambda\;|\;\lambda m=0\;\;\forall m\in M\}.$
Lemma 4.21.
We get that $\textup{ann}(M,\delta)$ is a differential graded twosided ideal of $(\Lambda,d)$.
Proof. Indeed, $\textup{ann}(M,\delta)$ is a twosided ideal by general ring theory. Further, if $\lambda\in\textup{ann}(M,\delta)$, then $\lambda m=0$ for any $m\in M$, and hence $0=\delta(\lambda m)=d(\lambda)m\pm\lambda\delta(m)=d(\lambda)m$ since $\delta(m)\in M$ and $\lambda\in\textup{ann}(M,\delta)$. This show that $d(\lambda)m=0$, and hence $d(\lambda)\in\textup{ann}(M,\delta)$.
Further, note that we always have $\textup{ann}(M,\delta)=\textup{ann}((M,\delta)[1])$.
Lemma 4.22.
Let $(\Lambda,d)$ be a differential graded algebra and let $(S,\delta)$ be a simple differential graded module. Then, $\textup{ann}(S,\delta)$ is a maximal differential graded twosided ideal and any simple $(\Lambda,d)/\textup{ann}(S,\delta)$ module is isomorphic to $(S,\delta)$.
Proof. We claim that $\textup{ann}(S,\delta)$ is a maximal twosided differential graded ideal. Suppose there is a larger twosided differential graded ideal $(X,d)$ of $(\Lambda,d)$ such that
$\textup{ann}(S,\delta)\subseteq(X,d)\subsetneq(\Lambda,d)$
But then we have an epimorphism
$(\Lambda/\textup{ann}(S,\delta),\overline{d})\longrightarrow\kern-8.00003pt\rightarrow(\Lambda/X,\overline{d}).$
Let $(T,\partial)$ be a simple $(\Lambda/\textup{ann}(S,\delta),\overline{d})$-module. Considering by inflation $(T,\partial)$ as a simple $(\Lambda,d)$-module, $\textup{ann}(T,\partial){\supseteq}\textup{ann}(S,\delta)$. Hence any maximal differential graded left ideal of $(\Lambda,d)$ containing $\textup{ann}(S,\delta)$ is a maximal differential graded left ideal of $(\Lambda,d)$ containing $\textup{ann}(T,\partial)$. But then, taking a non zero $x\in\ker(\delta)$ and a non zero $y\in\ker(\partial)$ the map
$S$$\longrightarrow$ $T$
$\lambda\cdot x$ $\mapsto$ $\lambda\cdot y$
defines a surjective homomorphism of differential graded $\Lambda$-modules. Since $(S,\delta)$ is simple, this map is an isomorphism.
Hence both algebras $(\Lambda,d)/\textup{ann}(S,\delta)$ and $(\Lambda,d)/(X,d)$ allow only one simple differential graded module, namely $(S,\delta)$. By definition $X$ acts trivially on $(S,\delta)$ and hence $X\subseteq\textup{ann}(S,\delta)$. We already assumed that $\textup{ann}(S,\delta)\subseteq(X,d)$, which shows that $X=\textup{ann}(S,\delta)$ is maximal.
Definition 4.23.
Let $(\Lambda,d)$ be a differential graded algebra.
  • Then the differential graded left radical $\textup{dgrad}_{\ell}(\Lambda,d)$ is the intersection of all maximal differential graded left ideals of $(\Lambda,d)$.
  • Likewise the differential graded right radical $\textup{dgrad}_{r}(\Lambda,d)$ is the intersection of all maximal differential graded right ideals of $(\Lambda,d)$.
  • The differential graded twosided radical $\textup{dgrad}_{2}(\Lambda,d)$ is the intersection of all of the annihilators of dg-simple differential graded left modules.
Proposition 4.24.
Let $(\Lambda,d)$ be a differential graded algebra. Then $\textup{dgrad}_{2}(\Lambda,d)$ is a twosided differential graded ideal. Suppose that $(\Lambda,d)$ is artinian and Noetherian as a differential graded left module. Then $\textup{dgrad}_{2}(\Lambda,d)$ is the smallest twosided differential graded ideal $I$ such that $(\Lambda/I,\overline{d})$ is a direct product of simple differential graded algebras. As a consequence, if $(\Lambda,d)$ is artinian as differential graded right module as well, then the intersection of the annihilators of all simple differential graded left modules coincides with the intersection of all differential graded simple right modules.
Proof. The fact that $\textup{dgrad}_{2}(\Lambda,d)$ is a twosided dg-ideal follows from Lemma 4.22. As $(\Lambda,d)$ is dg-Noetherian, there are maximal differential graded left ideals, and hence also differential graded simple $(\Lambda,d)$ dg-modules. To simplify notation, put
$J:=\bigcap_{(S,\delta)\textup{ simple dg module}}\textup{ann}(S,\delta).$
We shall prove that $(\Lambda/J,\overline{d})$ is a direct product of simple differential graded algebras.
We first remark that any simple differential graded $(\Lambda/J,\overline{d})$-module is a simple differential graded $(\Lambda,d)$-module as well. But also the converse holds, since $J$ is precisely the intersection of all annihilators of simples, so that the simple differential graded $(\Lambda,d)$-modules coincide with the simple differential graded $(\Lambda/J,\overline{d})$-modules.
We may now replace $(\Lambda,d)$ by $(\Lambda/J,\overline{d})$ and hence we shall assume now that $J=0$. By Lemma 4.21 we see that $\textup{ann}(S,\delta)$ is a twosided differential graded $(\Lambda,d)$-ideal. By Lemma 4.22 we have that $\textup{ann}(S,\delta)$ is a maximal twosided differential graded ideal. But then we have that $(\Lambda/\textup{ann}(S,\delta),\overline{d})$ is a simple differential graded algebra. Since $J=0$, we get that
$\Lambda=\prod_{(S,\delta)\textup{ simple dg module}}\Lambda/\textup{ann}(S,\delta).$
By Lemma 4.22 each of the twosided ideals $\textup{ann}(S,\delta)$ is maximal twosided differential graded and by Corollary 4.14 each $\Lambda/\textup{ann}(S,\delta)$ is a simple differential graded algebra. Conversely, if $(M,d)$ is a maximal differential graded twosided ideal, then again by Corollary 4.14 $(\Lambda/M,\overline{d})$ is differential graded simple. Since $(\Lambda,d)$ is artinian, so is $(\Lambda/M,\overline{d})$, and hence there are simple differential graded $(\Lambda/M,\overline{d})$-modules. Let $(S,\delta)$ be a simple differential graded $(\Lambda/M,\overline{d})$-module. Then $\textup{ann}(S,\delta)$, taken in $(\Lambda,d)$, is a maximal twosided differential graded ideal of $\Lambda$ containing $M$. Hence $M=\textup{ann}(S,\delta)$ and therefore $\textup{dgrad}_{2}(\Lambda,d)$ is the smallest twosided ideal with the required property.
Further this property is left-right invariant, and hence we obtain the same $\textup{dgrad}_{2}(\Lambda,d)$ in case we consider annihilators of simple differential graded right modules. The last statement now follows as well.
Proposition 4.25.
Let $(\Lambda,d)$ be a differential graded algebra, and suppose that $(\Lambda,d)$ is artinian as differential graded left and as a differential graded right module. Then
$\textup{dgrad}_{\ell}(\Lambda,d)\cap\textup{dgrad}_{r}(\Lambda,d)\supseteq\textup{dgrad}_{2}(\Lambda,d).$
Proof. We shall prove that $\textup{dgrad}_{2}(\Lambda,d)\subseteq\textup{dgrad}_{\ell}(\Lambda,d)$. The analogous statement for $\textup{dgrad}_{r}(\Lambda,d)$ is symmetric. By Lemma 4.15 every simple differential graded $(\Lambda,d)$-module is isomorphic to one of the form $(\Lambda/M,\overline{d})$ for a maximal differential graded ideal $(M,d)$ of $(\Lambda,d)$. If $\lambda\in\textup{ann}(\Lambda/M,\overline{d})$, then we obtain $\lambda\cdot(\Lambda/M)=0$, and hence $\lambda\cdot\Lambda\subseteq M$. This implies $\lambda\in M$ since $1\in\Lambda$. Therefore $(M,d)\supseteq\textup{ann}(\Lambda/M,\overline{d})$ and we get that the intersection of all ideals of the form $\textup{ann}(S,\delta)$ for simple dg-modules $(S,\delta)$ is contained in the intersection of all maximal differential graded ideals. This is $\textup{dgrad}_{2}(\Lambda,d)$. We therefore get
$\bigcap_{(S,\delta)\textup{ simple dg module}}\textup{ann}(S,\delta)\subseteq\textup{dgrad}_{2}(\Lambda,d).$
This shows the proposition.
Remark 4.26.
The hypothesis that $(\Lambda,d)$ is dg-artinian is only used to show that there are dg-simple dg-left modules. Just like in the classical situation, an easy use of Zorn’s lemma shows that it is sufficient to ask that $(\Lambda,d)$ is finitely generated as dg-left module.
Example 4.27.
Let $K$ be a field. Recall from Example 3.5 that the graded endomorphism algebra $\left(\begin{array}[]{cc}K&K\\ K&K\end{array}\right)$ of the complex $\textup{cone}(\textup{id}_{K})$ is a dg-algebra. It is not difficult to verify that $\left(\begin{array}[]{cc}0&K\\ 0&K\end{array}\right)=:I$ is the only non trivial differential graded ideal of $A$. Hence $\textup{dgrad}_{\ell}(A)=I$ and likewise the first row is the only non trivial differential graded right ideal. Hence $\textup{dgrad}_{\ell}(A)\cap\textup{dgrad}_{r}(A)=\left(\begin{array}[]{cc}0&K\\ 0&0\end{array}\right)$ is not an ideal at all. Further it strictly contains $\textup{dgrad}_{2}(A)=0$.
Proposition 4.28.
Let $(\Lambda,d)$ be a differential graded algebra. Suppose that $(\Lambda,d)$ is Noetherian and artinian as differential graded left module over itself. Then $\Lambda/\textup{dgrad}_{\ell}(\Lambda)$ is a direct sum of dg-simple dg-modules over $(\Lambda,d)$. In particular, $\textup{dgrad}_{\ell}(\Lambda)=0$ if and only if any differential graded $(\Lambda,d)$-module is isomorphic to a direct sum of simple differential graded $(\Lambda,d)$-modules.
Proof. We shall show that there are only finitely many maximal differential graded left ideals $I_{1},\dots,I_{n}$ of $\Lambda$, such that $I_{j}\not\supseteq\bigcap_{i=1}^{j-1}I_{i}$. Take a maximal differential graded left ideal $(I_{1},d)$. If there is no other maximal differential graded left ideal $(I_{2},d)$, we are done. Else we consider $(I_{1}\cap I_{2},d)$. If there is no maximal differential graded left ideal $I_{3}$ with $(I_{3},d)\not\supseteq I_{1}\cap I_{2}$, we are done. Else we consider $(I_{1}\cap I_{2}\cap I_{3},d)$. Inductively we obtain a chain of strictly decreasing differential graded left ideals. Since $\Lambda$ is assumed to be artinian, this chain is finite. Hence for any set of maximal differential graded ideals there is a finite subset of maximal differential graded ideals $I_{1},\dots,I_{n}$ such that $\bigcap_{i=1}^{n}I_{i}$ is contained in any maximal differential graded left ideal of $(\Lambda,d)$. Hence
$I_{1}\cap\dots\cap I_{n}=\textup{dgrad}_{\ell}(\Lambda)=\bigcap_{I\textup{ maximal dg ideal}}I.$
Then there is a homomorphism
$\Lambda\stackrel{{\scriptstyle\rho}}{{\longrightarrow}}\Lambda/I_{1}\times\dots\times\Lambda/I_{n}$
with kernel $\textup{dgrad}_{\ell}(\Lambda)$. Hence $\rho$ induces a monomorphism
$\Lambda/\textup{dgrad}_{\ell}(\Lambda)\stackrel{{\scriptstyle\overline{\rho}}}{{\hookrightarrow}}\Lambda/I_{1}\times\dots\times\Lambda/I_{n}.$
Lemma 4.16 then shows that there is a differential graded module $M$ such that
$\Lambda/\textup{dgrad}_{\ell}(\Lambda)\oplus M\simeq\Lambda/I_{1}\times\dots\times\Lambda/I_{n}.$
Since $\Lambda$ is Noetherian and artinian, the Krull-Schmidt theorem applies and hence, renumbering if necessary, there is $k$ such that
$\Lambda/\textup{dgrad}_{\ell}(\Lambda)\simeq\Lambda/I_{1}\times\dots\times\Lambda/I_{k}\textup{ and }M\simeq\Lambda/I_{k+1}\times\dots\times\Lambda/I_{n}.$
However, by construction of the sequence of ideals $I_{1},\dots,I_{n}$ we have $k=n$. Hence $\overline{\rho}$ is an isomorphism.
If $\Lambda$ is a direct sum of simple differential graded left modules, $\Lambda=S_{1}\times\dots\times S_{n}$, and hence a maximal dg-ideal is $S_{1}\times\dots\times S_{i-1}\times S_{i+1}\times\dots S_{n}$, and therefore the intersection of maximal dg-ideals is 00. This shows the proposition.
Lemma 4.29.
(differential graded Nakayama’s Lemma) Let $(\Lambda,d)$ be a differential graded algebra. Let $M$ and $N$ be dg-Noetherian and dg-artinian differential graded $\Lambda$-modules such that $N\leq M$. Then $\textup{dgrad}_{2}(\Lambda)M=M$ implies $M=0$, and $N+\textup{dgrad}_{2}(\Lambda)M=M$ implies $M=N$.
Proof. If $M=0$, the statement is trivial. Else, the hypothesis on $M$ implies that there is a dg-simple dg-module, namely a dg-simple dg-submodule of $M$. We will show the first statement. So suppose $\textup{dgrad}_{2}(\Lambda)M=M$. If $M$ is a dg-simple differential graded module, then $\textup{dgrad}_{2}(\Lambda)M=0$ since $\textup{dgrad}_{2}(\Lambda)$ is the intersection of the annihilators of the simples, and hence $M=0$. As $M$ is dg-Noetherian and dg-artinian, $M$ has a dg-composition series. We shall prove the statement by induction on the composition length. Since $M$ is dg-artinian, it has a differential graded dg-simple submodule $S$, and by construction $\textup{dgrad}_{2}(\Lambda)S=0$. Now, if $M\neq 0$ the composition length of $M$ is strictly bigger than the composition length of $\textup{dgrad}_{2}(\Lambda)M$. This provides a contradiction.
As for the second statement, $N+\textup{dgrad}_{2}(\Lambda)M=M$ implies $\textup{dgrad}_{2}(\Lambda)(M/N)=M/N$. The first statement then implies $M/N=0$, and therefore $M=N$. Hence the result follows.
Remark 4.30.
  1. (1)
    If $\Lambda$ is artinian forgetting the grading and the differential, it is Noetherian and artinian by Hopkins’ theorem, and so any finitely generated $\Lambda$-module $M$ is Noetherian and artinian as well.
  2. (2)
    Nakayama’s lemma in ring theory only needs that $M$ is Noetherian. However, the standard proof [24, Lemma 1.6.5] uses principal ideals and an elementwise argument. This does not seem to be available in the differential graded setting.
Remark 4.31.
There is a parallel work on Nakayama’s Lemma for differential graded algebras by Goodbody [10], based on the work of Orlov [17]. These results are very different from ours, and in particular Goodbody uses that the algebra is finite dimensional over some field. I discovered both articles after having submitted this manuscript.
5. Differential Graded Orders
Let $R$ be a Dedekind domain with field of fractions $K$. Recall that $K$ is a flat $R$-module.
Definition 5.1.
A differential graded $R$-order (dg-$R$-order) is a differential graded $R$-algebra $(\Lambda,d)$ such that $\Lambda$ is $R$-free and such that $K\otimes_{R}A$ is semisimple artinian. A differential graded $R$-order is a proper differential graded $R$-order if in addition $H(A)$ is $R$-free,
Definition 5.2.
Let $(\Lambda,d)$ be a dg-$R$-order. A differential graded $\Lambda$-lattice (dg-$R$-lattice) is a differential graded $\Lambda$-module $(L,d_{L})$ such that $L$ is $R$-free. A dg-lattice $(L,d_{L})$ is a proper dg-lattice if in addition $H(L)$ is $R$-free.
Remark 5.3.
Note that if $(\Lambda,d)$ is a dg-order, then a dg-lattice $(L,d_{L})$ can be proper or not. If $(\Lambda,d)$ is a proper dg-order, then a dg-lattice $(L,d_{L})$ can be proper or not.
Remark 5.4.
Let $R$ be a Dedekind domain with field of fractions $K$. Following Theorem 2.6 we may as well consider differential graded $K$-algebras $(A,d)$ which are exact and $\ker(d)$ semisimple. Then we may define an $R$-order as a differential graded $R$-subalgebra $(\Lambda,d)$ projective as an $R$-module with $K\otimes_{R}\Lambda=A$. This will give a different definition, implying that the homology is necessarily $R$-torsion. Further, the fact that the algebra $A$ is very general does not seem to allow a rich theory. Moreover, as a classical order is not a dg-order with differential 00 in this case, this alternative theory will not generalize the classical theory of lattices over orders, and in particular it is unclear how one might recover in a profitable way complexes of lattices over a classical order in any way as differential graded module over a dg-order in this alternative definition.
Example 5.5.
We recall some examples of classical orders and explain differential graded structures on them.
  1. (1)
    For $R$ a complete discrete valuation ring containing ^$\widehat{\mathbb{Z}}_{3}$ the $3$-adic integers, following [25, Example 1.2.43.4] the group ring $R{\mathfrak{S}}_{3}$ of the symmetric group of order $6$ is isomorphic to
    $\{(d_{1},\left(\begin{array}[]{cc}a_{2}&b_{2}\\ c_{2}&d_{2}\end{array}\right),a_{3})\in R\times Mat_{2\times 2}(R)\times R\;|\;a_{2}-d_{1}\in 3R,c_{2}\in 3R,a_{3}-d_{2}\in 3R\}.$
    This algebra is differential graded with the grading given by $deg(a_{i})=deg(d_{i})=0$, $deg(c_{2})=-1$ and $deg(b_{2})=1$ for all $i$, using the designation of the variables as in the above definition. The differential is given as in Example 3.5 by $d_{x}$ for any fixed $x\in R$. More precisely
    $d_{x}((d_{1},\left(\begin{array}[]{cc}a_{2}&b_{2}\\ c_{2}&d_{2}\end{array}\right),a_{3}))=(0,\left(\begin{array}[]{cc}xc_{2}&x(d_{2}-a_{2})\\ 0&xc_{2}\end{array}\right),0).$
    We can even consider suborders, such as those given by asking in addition that $a_{2}-d_{2}\in 3R$.
  2. (2)
    Fix a prime $p>2$. We know that the principal block of $R{\mathfrak{S}}_{p}$, for $R$ being a complete discrete valuation ring containing $\widehat{Z}_{p}$, the $p$-adic integers and ${\mathfrak{S}}_{p}$ the symmetric group of degree $p$, is Morita equivalent to
    $\{(d_{1},\prod_{i=2}^{p-1}\left(\begin{array}[]{cc}a_{i}&b_{i}\\ c_{i}&d_{i}\end{array}\right),a_{p})\in R\times\prod_{i=2}^{p-1}Mat_{2\times 2}(R)\times R\;|\;a_{j+1}-d_{j}\in pR,c_{j}\in pR,\forall j\}.$
    Again, following Example 3.5, for every of the $2\times 2$ matrix algebras in the product we may impose a differential graded algebra structure with differential $d_{x_{i}}$ by choosing $x_{2},\dots,x_{p-1}\in R$.
  3. (3)
    Fix a prime $p>2$ and let $R$ be a complete discrete valuation ring containing $\widehat{Z}_{p}$, the $p$-adic integers. By a result due to König [14] the principal block of the Schur algebra $S_{R}(p,p)=End_{R{\mathfrak{S}}_{p}}((R^{p})^{\otimes p})$ with parameters $(p,p)$ is known to be Morita equivalent to an algebra of similar shape.
    $\{(\prod_{i=1}^{p-1}\left(\begin{array}[]{cc}a_{i}&b_{i}\\ c_{i}&d_{i}\end{array}\right),a_{p})\in\prod_{i=1}^{p-1}Mat_{2\times 2}(R)\times R\;|\;a_{j+1}-d_{j}\in pR,c_{j}\in pR,\forall j\}.$
    The same statement on differential graded structures hold.
  4. (4)
    Fix a prime $p$. By [23] ([6, 7] for the case $p=2$ and $p=3$) the category ${\mathcal{F}}^{p}_{{\mathbb{Z}}}$ of polynomial functors of degree at most $p$ from finitely generated free abelian groups to finitely generated $\widehat{Z}_{p}$-modules is equivalent to a module category, having a ’principal block’ and a number of trivial direct factors. The principal block is Morita equivalent to
    $\{(d_{0},\prod_{i=1}^{p-1}\left(\begin{array}[]{cc}a_{i}&b_{i}\\ c_{i}&d_{i}\end{array}\right),a_{p})\in\widehat{\mathbb{Z}}_{p}\times\prod_{i=1}^{p-1}Mat_{2\times 2}(\widehat{\mathbb{Z}}_{p})\times\widehat{\mathbb{Z}}_{p}\;|\;a_{j+1}-d_{j}\in p\widehat{\mathbb{Z}}_{p},c_{j}\in p\widehat{\mathbb{Z}}_{p},\forall j\}.$
    Again, following Example 3.5, for every of the $2\times 2$ matrix algebras in the product we may impose a differential graded algebra structure with differential $d_{x_{i}}$ by choosing $x_{1},\dots,x_{p-1}\in p\widehat{\mathbb{Z}}_{p}$.
Remark 5.6.
Le Bruyn et al studied in [15, Section II.4] graded orders over Krull domains in a very general setting. They define for a Krull domain $R$ with field of fractions $K$ and a central simple $K$-algebra $A$ a subring $\Lambda$ with $R\subset\Lambda\subset A$ an $R$-order if each element $a$ of $\Lambda$ is integral over $R$, and if $K\Lambda=A$. If $a$ is integral over $R$, then for the reduced trace $Tr:A\rightarrow K$ one has $Tr(a)\in R$ and $Tr$ induces an isomorphism $A\rightarrow Hom_{K}(A,K)$ by $A\ni a\mapsto(A\ni b\mapsto Tr(ab)\in K)\in Hom_{K}(A,K)$.
Recall in this context the classical result
Theorem 5.7.
(cf. e.g. [19], [25, Theorem 8.3.7]) Let $R$ be a Dedekind domain of characteristic 00 with field of fractions $K$ and let $\Lambda$ be an $R$-subalgebra of the semisimple $K$-algebra $A$, containing a $K$-basis of $A$. Then $\Lambda$ is an $R$-order if and only if every element of $\Lambda$ is integral over $R$.
Hence the definition of [15, Section II.4] coincides with the classical definition of an $R$-order in case of a Dedekind domain $R$.
Recall that a dg-module $(V,\delta)$ with $\delta(V_{i})\subseteq V_{i+1}$ for all $i$ is right bounded if there is $n_{0}\in{\mathbb{N}}$ such that $V_{n}=0$ for all $n>n_{0}$. Analogously we define left bounded dg-modules. A dg-module is bounded if it is at once left and right bounded. Note that for a differential graded $R$-order $(\Lambda,d)$ in a finite dimensional semisimple differential graded $K$-algebra $(A,d)$ the lattice $(\Lambda,d)$ is always bounded. Indeed, since $(A,d)$ is finite dimensional, $(A,d)$ is bounded, and hence so is $(\Lambda,d)$.
Lemma 5.8.
Let $R$ be a Dedekind domain with field of fraction $K$. Let $(\Lambda,d)$ be a differential graded $R$-order in a finite dimensional differential graded $K$-algebra $(A,d)$ and let $(V,\delta)$ be finite dimensional differential graded (hence bounded) $(A,d)$-module. Then there is a differential graded $(\Lambda,d)$-lattice $(L,\delta)$ in $(V,d)$ such that $K\otimes_{R}L=V$.
Proof. Recall that classically for an $R$-order $\Lambda$ and a $K\Lambda$-module $V$ there is always a full $\Lambda$-lattice $L$ in $V$. Consider $\Lambda^{+}:=\bigoplus_{k\geq 1}\Lambda_{k}$ and $\Lambda_{0}^{+}:=\Lambda_{0}\oplus\Lambda^{+}$. Now, $\Lambda_{0}^{+}$ is a (differential graded) subalgebra of $\Lambda$ and $\Lambda^{+}$ is a dg-ideal of $\Lambda_{0}^{+}$.
Let us perform the differential graded construction. We construct $(L,\delta)$ by downward induction on the degree. Let $V_{k}=0\neq V_{n-1}$ for $k\geq n$ and take any $\Lambda_{0}^{+}$-lattice in $V_{n-1}$. Its existence is a well-known classical property for lattices. Note that $\Lambda^{+}$ acts as 00. Hence this is automatically a $(\Lambda_{0}^{+},d)$-lattice since $\delta(V_{n-1})=0$ and hence the Leibniz formula automatically holds.
Let $\widetilde{L}_{n-2}:=\delta^{-1}(L_{n-1})\subseteq V_{n-2}$ and choose a full $\Lambda_{0}$-lattice $\widehat{L}_{n-2}$ in $\widetilde{L}_{n-2}$. Then choose an $r\in R\setminus\{0\}$ such that $r\cdot\Lambda^{+}\cdot\widehat{L}_{n-2}\subseteq L_{n-1}$. This is possible since $\Lambda^{+}\cdot\widehat{L}_{n-2}$ is a $\Lambda^{+}_{0}$-lattice. Put $L_{n-2}:=r\cdot\widehat{L}_{n-2}$. Then by construction $\delta(L_{n-2})\subseteq L_{n-1}$ and $L_{n-2}\oplus L_{n-1}$ is a $\Lambda^{+}_{0}$-lattice, and moreover the Leibniz formula holds since it holds for the $(A,d)$-action on $(V,\delta)$. We suppose having constructed $L_{k},L_{k+1},\dots,L_{n}$. Then let $\delta^{-1}(L_{k})=:\widetilde{L}_{k-1}$ and choose a $\Lambda_{0}$-lattice $\widehat{L}_{k-1}$ in $\widetilde{L}_{k-1}$. Then there is again an $r\in R\setminus\{0\}$ such that
$r\cdot\Lambda^{+}\cdot\widehat{L}_{k-1}\subseteq\bigoplus_{\ell=k}^{n}L_{\ell}.$
Put $L_{k-1}:=r\cdot\widehat{L}_{k-1}$. Again, $\bigoplus_{\ell=k-1}^{n}L_{\ell}$ this is a full differential graded $\Lambda^{+}_{0}$-lattice. Since $(V,\delta)$ is bounded, after a finite number of steps we constructed a Noetherian differential graded $\Lambda^{+}_{0}$-lattice $(L,\delta)$ as $\bigoplus_{k\in{\mathbb{Z}}}L_{k}$.
Now, since $\Lambda$ is Noetherian as well, $\Lambda\cdot L$ is Noetherian again, and since it is a submodule of $V$, it is $R$-torsion free. Hence $\Lambda\cdot L$ is a lattice in $(V,\delta)$. Moreover, the Leibniz formula holds. Since
$\delta(\lambda\cdot x)=d(\lambda)\cdot x\pm\lambda\cdot\delta(x)\in\Lambda\cdot L$
for homogeneous elements $\lambda\in\Lambda$ and $x\in L$. This shows that
$\delta(\Lambda\cdot L)\subseteq\Lambda\cdot L$
is a full differential graded $(\Lambda,d)$-lattice in $(V,d)$.
Remark 5.9.
Note that we do not really need $\Lambda_{0}^{+}$. It is possible to first find a full and finitely generated $R$-submodule $\check{L}$ of $V$, stable under $\delta$, and then consider $\Lambda\cdot\check{L}$. Noetherianity and the last argument of the above proof then provides the result. However, the above proof of Lemma 5.8 gives a more direct construction for coconnective dg-algebras. A similar construction can be given for connective dg-algebras.
In the classical theory of orders the following result is a main tool. We shall need to transpose it to the differential graded situation.
Proposition 5.10.
Let $R$ be a Dedekind domain with field of fractions $K$. For each $\wp\in\textup{Spec}(R)$ and any $R$-module $M$ denote $M_{\wp}:=R_{\wp}\otimes_{R}M$ and $\textup{id}_{R_{\wp}}\otimes d:=d_{\wp}$ for any map $d$, in particular the differential.
  1. (1)
    If $(\Lambda,d_{\Lambda})$ is a (resp. proper) dg-$R$-order in the semisimple dg-$K$-algebra $(A,d_{A})$, then $(\Lambda_{\wp},1_{R_{\wp}}\otimes d_{\Lambda})$ is a (resp. proper) dg-$R_{\wp}$-order in the semisimple dg-$K$-algebra $(A,d_{A})$.
  2. (2)
    If $(L,d_{L})$ is a (resp. proper) dg-$R$-lattice, then $(L_{\wp},(d_{L})_{\wp})$ is a (resp. proper) dg-$(\Lambda_{\wp},(d_{\Lambda})_{\wp})$-lattice.
  3. (3)
    For any (resp. proper) dg-$(\Lambda,d_{\Lambda})$-lattice $(L,d_{L})$ we have
    $L=\bigcap_{\wp\in\textup{Spec}(R)}(L_{\wp},(d_{L})_{\wp})$
    where the intersection is taken inside $K\otimes_{R}L$.
  4. (4)
    Fix a dg-$(A,d_{A})$-module $(V,d_{V})$. For each $\wp\in\textup{Spec}(R)$ fix (resp. proper) $(\Lambda_{\wp},(d_{\Lambda})_{\wp})$-lattices $(M(\wp),(d_{M}(\wp)))$ such that $(KM(\wp),(Kd_{M}(\wp)))=(V,d_{V})$ for all $\wp\in\textup{Spec}(R)$. Suppose moreover that there is a (resp. proper) dg-$(\Lambda,d_{\Lambda})$-lattice $(N,d_{N})$ such that $(N_{\wp},(d_{N})_{\wp})=(M(\wp),d_{M}(\wp))$ for all but a finite number of $\wp\in\textup{Spec}(R)$. Then there is a (resp. proper) dg-$(\Lambda,d_{\Lambda})$-lattice $(L,d_{L})$ with $(L_{\wp},(d_{L})_{\wp})=(M(\wp),d_{M}(\wp))$ for all $\wp\in\textup{Spec}(R)$.
  5. (5)
    The analogous statements hold replacing the localisation by the completion.
Proof. Item (1) follows from the classical non dg-statement (cf e.g. [25, Proposition 8.1.14]).
For item (2) we need to see that $H(L_{\wp},(d_{L})_{\wp})$ is $R_{\wp}$-torsion free if $H(L,d_{L})$ is $R$-torsion free. But this follows from the fact that localisation is flat (cf e.g. [25, Lemma 6.5.7]).
Item (3) is again a direct consequence of the classical non dg-statement (cf e.g. [25, Proposition 8.1.14]).
As for item (4) we first get a lattice $L$ again from the classical situation (cf e.g. [25, Proposition 8.1.14]). The differential is fixed as the restriction of the differential $d_{V}$ on $L$. We need to verify that $d_{V}(L)\subseteq L$. But this is true at every prime, i.e.
$d_{V}(L_{\wp})=d_{M}(\wp)(M(\wp))\subseteq M(\wp)=L_{\wp}.$
Hence by the non dg-version of item (3) we have $d_{V}(L)\subseteq L$.
As for item (5) we first see that the non dg-version is again classical (cf e.g. [25, Proposition 8.1.14]). The analogous of the second statement first uses the localisation, then the uniqueness and existence of the differential follows by continuity. The third item is clear by the localisation case. As for the fourth item we first restrict from the completion to $A$. Then, in a second step we use continuity again.
An important tool in case of lattices for orders is the conductor. Recall that given a Dedekind domain $R$ with field of fractions $K$ and a finite dimensional semisimple $K$-algebra $A$, then for an $R$-lattice $L$ in $A$ with $KL=A$ we define
${\mathcal{O}}_{\ell}(L):=\{\lambda\in A\;|\;\lambda L\subseteq L\}\text{ and }{\mathcal{O}}_{r}(L):=\{\lambda\in A\;|\;L\lambda\subseteq L\}.$
Then by [25, Lemma 8.3.14] we get that ${\mathcal{O}}_{\ell}(L)$ and ${\mathcal{O}}_{r}(L)$ are $R$-orders in $A$.
We can show a dg-version of this lemma.
Lemma 5.11.
Let $R$ be a Dedekind domain with field of fractions $K$ and let $(\Lambda,d)$ be a differential graded $R$-order in the semisimple differential graded $K$-algebra $(A,d)$. Let $(L,d_{L})$ be a differential graded $R$-lattice in $(A,d)$ such that $K\cdot(L,d_{L})=(A,d)$. Then $({\mathcal{O}}_{\ell}(L),d|_{{\mathcal{O}}_{\ell}(L)})$ is a differential graded $R$-order in $(A,d)$. Similar statements hold for ${\mathcal{O}}_{r}(L)$.
Proof. By symmetry it is enough to consider the case $({\mathcal{O}}_{\ell}(L),d|_{{\mathcal{O}}_{\ell}(L)})$. By [25, Lemma 8.3.14] we see that ${\mathcal{O}}_{\ell}(L)$ is an $R$-order in $A$. We need to see that it is differential graded. Let $\lambda\in{\mathcal{O}}_{\ell}(L)$ be homogeneous. Then $\lambda L\subseteq L$. We need to see that $d(\lambda)L\subseteq L$. Since $L\subseteq A$, and since $(A,d)$ is a differential graded algebra, we have for any $x\in L$ and homogeneous $\lambda\in{\mathcal{O}}_{\ell}(L)$
$d(\lambda x)=d(\lambda)x+(-1)^{|\lambda|}\lambda d(x)$
Hence
$d(\lambda)x=d(\lambda x)-(-1)^{|\lambda|}\lambda d(x)$
Since $\lambda\in{\mathcal{O}}_{\ell}(L)$ and since $d(L)\subseteq L$, using that $(L,d)$ is a dg-lattice, we have $\lambda d(x)\in L$. Since $\lambda\in{\mathcal{O}}_{\ell}(L)$, and since $x\in L$, we get $\lambda x\in L$, and hence $d(\lambda x)\in L$ again. The differential on ${\mathcal{O}}_{\ell}(L)$ is the restriction of the differential $d$ of $A$ to ${\mathcal{O}}_{\ell}(L)$, and hence the defining equation on products holds. Hence $({\mathcal{O}}_{\ell}(L),d|_{{\mathcal{O}}_{\ell}(L)})$ is a dg-order in $(A,d)$.
Remark 5.12.
If $(A,d)$ is semisimple as a dg-algebra, then by [1] we have that $H(A,d)=0$ and therefore any order in $(A,d)$ is either exact or has $R$-torsion homology. In particular, in this case, if the conductor is not exact, $({\mathcal{O}}_{\ell}(L),d|_{{\mathcal{O}}_{\ell}(L)})$ cannot be a proper differential graded $R$-order in $(A,d)$, whatever the choice of a dg-lattice $L$ may be.
6. On maximal differential graded orders
Let $R$ be a Dedekind domain with field of fractions $K$. It is a classical fact that for any semisimple $K$-algebra $A$ and an $R$-order $\Lambda$ in $A$ there is a maximal $R$-order $\Gamma$ containing $\Lambda$. Maximal orders have many striking properties, and behave very much as the base ring $R$ (cf [19]). We shall consider maximal dg-orders.
Theorem 6.1.
Let $R$ be a Dedekind domain with field of fractions $K$ and let $(\Lambda,d)$ be a differential graded $R$-order in the semisimple finite dimensional differential graded $K$-algebra $(A,d)$. Then there is a maximal differential graded $R$-order $(\Gamma,d)$ containing $(\Lambda,d)$. If $(\Lambda,d)$ is a proper dg-order, then there is a maximal proper differential graded $R$-order $(\Gamma_{p},d)$ containing $(\Lambda,d)$.
Proof. The set of dg-orders in $(A,d)$ is partially ordered. We need to make precise the partial ordering we consider. Let $(\Lambda_{1},d_{1})$ and $(\Lambda_{2},d_{2})$ be dg-$R$-orders in a semisimple dg-$K$-algebra $(A,d)$. Then $(\Lambda_{1},d_{1})\leq(\Lambda_{2},d_{2})$ if $\Lambda_{1}\subseteq\Lambda_{2}$. Since $d_{1}=d|_{\Lambda_{1}}$ and $d_{2}=d|_{\Lambda_{2}}$, we get that then automatically $d_{1}=d_{2}|_{\Lambda_{1}}$.
Consider
${\mathcal{X}}:=\{(\Sigma,d)\;|\;(\Lambda,d)\leq(\Sigma,d)\;\textup{ and }(\Sigma,d)\textup{ is a dg-$R$-order in $(A,d)$}\}.$
Since $(\Lambda,d)\in{\mathcal{X}}$, this set ${\mathcal{X}}$ is not empty. Let ${\mathcal{Y}}$ be a totally ordered subset in ${\mathcal{X}}$ and put $\Gamma:=\bigcup_{(\Delta,d)\in{\mathcal{Y}}}\Delta$. It is graded since the embedding of one order in the other in ${\mathcal{Y}}$ preserves the grading, and hence so is the union. It allows a differential by $d|_{\Gamma}$ and if $\gamma\in\Gamma$, then $\gamma\in\Delta$ for some $\Delta\in{\mathcal{Y}}$, and therefore $d(\gamma)\in\Delta\subseteq\Gamma$. Moreover, since $\gamma\in\Delta$, and since $\Delta$ is an order, Theorem 5.7 shows that $\gamma$ is integral, and, using Theorem 5.7 again, $\Gamma$ is an order in $A$. Hence $(\Gamma,d)\in{\mathcal{X}}$ and $(\Gamma,d)$ dominates any element in $\mathcal{Y}$. By Zorn’s lemma there are maximal elements in $\mathcal{X}$. Any such element is a maximal differential graded order containing $(\Lambda,d)$.
Now consider
$\widehat{\mathcal{X}}:=\{(\Sigma,d)\;|\;(\Lambda,d)\leq(\Sigma,d)\;\textup{ and }(\Sigma,d)\textup{ is a proper dg-$R$-order in $(A,d)$ and $\Sigma$}\}$
and if $(\Lambda,d)$ is a proper dg-$R$-order in $(A,d)$, then $(\Lambda,d)\in\widehat{\mathcal{X}}$. Hence this set is not empty neither. The first steps are as above. Let $\widehat{\mathcal{Y}}$ be a totally ordered subset in $\widehat{\mathcal{X}}$ and put $\Gamma:=\bigcup_{(\Delta,d)\in\widehat{\mathcal{Y}}}\Delta$. It is graded since the embedding of one order in the other in ${\mathcal{Y}}$ preserves the grading, and hence so is the union. It allows a differential by $d|_{\Gamma}$ and if $\gamma\in\Gamma$, then $\gamma\in\Delta$ for some $\Delta\in\widehat{\mathcal{Y}}$. Hence $\gamma$ is integral, and therefore $\Gamma$ is an order in $A$. Since $R$ is Noetherian (cf [25, Lemma 7.5.3]), $\Gamma$ is Noetherian as well, and therefore $\Gamma$ actually in $\widehat{\mathcal{Y}}$. Therefore $(\Gamma,d)$ is a proper dg-order. By Zorn’s lemma, $\widehat{\mathcal{X}}$ contains maximal elements. This proves the statement.
Corollary 6.2.
Let $R$ be a Dedekind domain with field of fractions $K$, let $(A,d)$ be a finite dimensional semisimple differential graded $K$-algebra, and let $\Lambda$ be a maximal $R$-order in $A$ and suppose that $d(\Lambda)\subseteq\Lambda$. Then $(\Lambda,d|_{\Lambda})$ is a maximal differential graded $R$-order in $(A,d)$.
Proof. The hypotheses imply that $(\Lambda,d|_{\Lambda})$ is a differential $R$-graded order. If $(\Gamma,d|_{\Gamma})$ is a differential graded order containing $(\Lambda,d|_{\Lambda})$, then $\Gamma$ is an $R$-order containing $\Lambda$. Hence, since $\Lambda$ is a maximal order, $\Lambda=\Gamma$ and therefore $(\Lambda,d|_{\Lambda})$ is a maximal differential $R$-graded order. This shows the statement.
Remark 6.3.
Note that in the situation of Corollary 6.2 there is no reason why $H(\Lambda,d_{\Lambda})$ should be $R$-projective, and indeed $H(A,d)=0$, as in Example 3.5, is possible and implies this case. Recall from Remark 2.6 that orders in algebras which are semisimple as differential graded algebra are actually necessarily of this form. Let $(\Lambda,d)$ be an order in $(A,d)$ with $H(\Lambda,d)=0$. Then Aldrich and Garcia-Rocas show in [1, Theorem 4.7] that in this case the category of differential graded $(\Lambda,d)$-modules is equivalent with the category of graded $\ker(d)$-modules.
Example 6.4.
Let $R$ be an integral domain with field of fractions $K$. We consider the differential graded semisimple $K$-algebra $(Mat_{2\times 2}(K),d_{x})$ from Example 3.5. Recall that for each $x\in K$ there is a differential $d_{x}$ on $Mat_{2\times 2}(K)$ with the grading chosen in Example 3.5.
Choose $x=1$ for the moment. Then $Mat_{2\times 2}(R)$ is a maximal differential graded $R$-order, and actually a proper differential graded $R$-order since the homology of $(Mat_{2\times 2}(R),d_{1})$ is 00.
However, choosing an element $x\in R\setminus R^{\times}$, then again $Mat_{2\times 2}(R)$ is a maximal differential graded $R$-order. The homology in degree $1$ is $R/xR$, as is easily seen. The kernel of the differential in degree 00 is $C_{0}:=\{\left(\begin{array}[]{cc}a&0\\ 0&a\end{array}\right)\;|\;a\in R\}$ and the image if the differential from degree $-1$ is $xC_{0}$. Clearly, the differential in degree $-1$ is injective. Hence $H_{*}(Mat_{2\times 2}(R),d_{x})\simeq R/xR\oplus R/xR$, where the first copy is in degree 00 and the second copy is in degree $1$. Therefore $H_{*}(Mat_{2\times 2}(R),d_{x})\simeq(R/xR)[\epsilon]/\langle\epsilon^{2}\rangle$ where $\epsilon$ is an element in degree $1$.
Note that $Mat_{2\times 2}(R)$ is hereditary, as an order, whereas, as soon as $x\not\in R^{\times}$, the homology algebra $H_{*}(Mat_{2\times 2}(R),d_{x})$ is not, even of infinite global dimension.
Example 6.5.
Let $R$ be an integral domain with field of fractions $K$. We consider again the differential graded semisimple $K$-algebra $(Mat_{2\times 2}(K),d_{x})$ from Example 3.5. If $R={\mathbb{Z}}$ and $x=\frac{1}{2}$, then for $\Lambda=Mat_{2\times 2}(R)$ we obtain $d_{x}(\Lambda)$ is not a subset of $\Lambda$. Actually,
$d_{\frac{1}{2}}(Mat_{2\times 2}({\mathbb{Z}}))=\frac{1}{2}{\mathbb{Z}}\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)+\frac{1}{2}{\mathbb{Z}}\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right).$
Since
$\left(\begin{array}[]{cc}\frac{1}{2}&0\\ 0&\frac{1}{2}\end{array}\right)\cdot\left(\begin{array}[]{cc}0&\frac{1}{2}\\ 0&0\end{array}\right)=\left(\begin{array}[]{cc}0&\frac{1}{4}\\ 0&0\end{array}\right)$
and iterating, we have that for each integer $n$ the element $\left(\begin{array}[]{cc}0&\frac{1}{2^{n}}\\ 0&0\end{array}\right)$ is in the ring generated by $d_{\frac{1}{2}}(Mat_{2\times 2}({\mathbb{Z}}))$. Hence, there is no differential graded order $(\Gamma,d|_{\Gamma})$ such that $Mat_{2\times 2}(K)\subseteq\Gamma$. However, for any $x\in K\setminus\{0\}$ the set
$\Lambda:=\left(\begin{array}[]{cc}R&xR\\ \frac{1}{x}R&R\end{array}\right)$
is a subring of $Mat_{2\times 2}(K)$ and is stable under $d_{x}$. Further, if $R$ is a Dedekind domain, then $xR$ is projective, and hence $\Lambda$ is an $R$-order if $R$ is a Dedekind domain. We observe that
$\left(\begin{array}[]{cc}\frac{1}{x}&0\\ 0&1\end{array}\right)\cdot\Lambda\cdot\left(\begin{array}[]{cc}x&0\\ 0&1\end{array}\right)=Mat_{2\times 2}(R)$
and hence $\Lambda$ is conjugate to the maximal order $Mat_{2\times 2}(R)$. Therefore, using Corollary 6.2, $(\Lambda,d_{x})$ is a maximal differential graded order. We get $H_{*}(\Lambda,d_{x})=0$, and hence this is actually a proper differential graded $R$-order.
Question 6.6.
If $(\Lambda,d)$ is a maximal differential graded order, can one show that $\Lambda$ is a maximal order.
Question 6.7.
If $(\Lambda,d)$ is a maximal proper differential graded order, can we show that $H(\Lambda,d)$ is hereditary?
7. Class groups of differential graded orders
Recall that a differential graded $R$-order $(\Lambda,d)$ in a semisimple differential graded $K$-algebra $(A,d)$ is at first an $R$-order $\Lambda$ in a semismple $K$-algebra $A$. The locally free class group of an order proved to be a useful invariant in integral representation theory. We want to provide a definition and first properties of a locally free class group of differential graded orders.
We first need to elaborate on what we mean by a free differential graded module. Let $(\Lambda,d)$ be a differential graded algebra. A differential graded $(\Lambda,d)$-module $(L,\delta)$ is free of rank $n\in{\mathbb{N}}$ if $(L,\delta)\simeq(\Lambda,d)^{n}$ as differential graded $(\Lambda,d)$-modules.
Since $d(1)=0$ in $\Lambda$, we get that for an isomorphism $\varphi:(\Lambda,d)^{n}\longrightarrow(L,\delta)$ we have $\varphi(0,\dots,0,1,0,\dots,0)\in\ker(\delta)$, where $1$ is in position $i$, for each position $i$. However, for any $z\in\ker(\delta)$ we get that $\varphi(\lambda):=\lambda z$ defines a homomorphism $(\Lambda,d)\longrightarrow(L,\delta)$.
Lemma 7.1.
Let $R$ be a Dedekind domain with field of fractions $K$. Let $(\Lambda,d)$ be a differential graded algebra, and denote $(A,d):=(K\otimes_{R}\Lambda,\textup{id}_{K}\otimes d)$. Then the set of free differential graded $(\Lambda,d)$-ideals is in bijection with the group of left regular elements in $\ker(d)\subseteq A$ modulo the subgroup $(\ker(d)\cap\Lambda^{\times})$.
Proof. We now consider dg-ideals $(I,d)$ of $(\Lambda,d)$. Such an ideal is free if there is an isomorphism $\varphi:(\Lambda,d)\longrightarrow(I,d)$ and this is equivalent with the choice of a $z\in(I\cap\ker(d))$ such that $I=\Lambda z$ and such that $\lambda z=0\Rightarrow\lambda=0$. We hence obtain that any left regular generator of $I$ in the cycles gives an isomorphism, and hence there is a surjective map from left regular elements in the cycles to the set of rank one free dg $(\Lambda,d)$ ideals. Two such left regular elements $z_{1},z_{2}$ of $I$ in the cycles give the same ideal if and only if $\Lambda z_{1}=\Lambda z_{2}$. This is equivalent with $z_{1}=\lambda_{2}z_{2}$ and $z_{2}=\lambda_{1}z_{1}$ for units $\lambda_{1},\lambda_{2}\in\Lambda$. However, $z_{1},z_{2}\in\ker(d)$ shows that
$0=d(\lambda_{2}z_{2})=d(\lambda_{2})z_{2}\pm\lambda_{2}d(z_{2})=d(\lambda_{2})z_{2}$
and the fact that $z_{2}$ is regular gives that $\lambda_{2}\in\ker(d)$. Likewise $\lambda_{1}\in\ker(d)$.
Conversely, if $\lambda\in\ker(d)$, then $\Lambda z=\Lambda\lambda z$ as differential graded ideal. Indeed, if $\lambda\in\ker(d)$ and $z\in\ker(d)$, then
$d(\mu\lambda z)=d(\mu)\lambda z+(-1)^{|\mu|}\mu d(\lambda)z+(-1)^{|\mu\lambda|}\mu\lambda d(z)=d(\mu)\lambda z$
and hence mapping $\mu z$ to $\mu\lambda z$ for any $\mu\in\Lambda$ is an isomorphism of differential graded ideals.
We shall need to study invertible elements in the subring of cycles.
Lemma 7.2.
Let $(\Lambda,d)$ be a differential graded algebra. Then $\ker(d)\cap\Lambda^{\times}=\ker(d)^{\times}$.
Proof. We have seen in Corollary 1.3 that $\ker(d)$ is a graded subalgebra. Trivially $\ker(d)\cap\Lambda^{\times}\supseteq\ker(d)^{\times}$. Further, if $u$ is invertible in $\Lambda$, and if $u\in\ker(d)$, then also $u^{-1}\in\ker(d)$. Indeed, $0=d(1)=d(uu^{-1})=d(u)u^{-1}\pm ud(u^{-1})$ and hence $d(u^{-1})=0$ since $u$ is invertible. Hence $\ker(d)\cap\Lambda^{\times}\subseteq\ker(d)^{\times}$.
Let $R$ be a Dedekind domain with field of fractions $K$. Suppose now that $(A,d)$ is a finite dimensional differential graded $K$-algebra and suppose that any left regular element of $A$ is invertible. By Wedderburn’s theorem, in a finite dimensional semisimple $K$-algebra an element $u$ is left regular if and only if it is represented by a tuple of non singular matrices, whence an invertible element. Let $(\Lambda,d)$ be a differential graded subalgebra of $(A,d)$ such that $\Lambda$ is $R$-projective and $K\otimes_{R}\Lambda=A$. As in the classical case we need to work with two versions, the localised version and the completed version. Let $X_{\wp}$ be the localisation at a prime Weierstrass-$\wp$, and denote by $\widehat{X}_{\wp}$ the completion at the prime Weierstrass-$\wp$. Hence we denote
$J(A,d):=\{(u(\wp))_{\wp\in Spec(R)}\in\prod_{\wp\in Spec(R)}\left(A_{\wp}^{\times}\cap\ker(d)\right)\;|\;u(\wp)\in\Lambda_{\wp}^{\times}\text{ for almostall $\wp\in Spec(R)$}\}.$
Note that this definition does not depend on the choice of $\Lambda$. Define its subgroup
$U(\Lambda,d):=\prod_{\wp\in Spec(R)}(\Lambda_{\wp}^{\times}\cap\ker(d)),$
which does depend on $\Lambda$. Likewise
$\widehat{J}(A,d):=\{(u(\wp))_{\wp\in Spec(R)}\in\prod_{\wp\in Spec(R)}\left(\widehat{A}_{\wp}^{\times}\cap\ker(d)\right)\;|\;u(\wp)\in\widehat{\Lambda}_{\wp}^{\times}\text{ for almost all $\wp\in Spec(R)$}\}.$
and its subgroup
$\widehat{U}(\Lambda,d):=\prod_{\wp\in Spec(R)}(\widehat{\Lambda}_{\wp}^{\times}\cap\ker(d)).$
Then the set of left classes
$U(\Lambda,d)\backslash J(A,d)$
is in bijection with the set of locally free differential graded fractional ideals of $(\Lambda,d)$ and
$\widehat{U}(\Lambda,d)\backslash\widehat{J}(A,d)$
is in bijection with the set of completion locally free differential graded fractional ideals of $(\Lambda,d)$.
We associate to a representative $\alpha=(\alpha(\wp))_{\wp\in Spec(R)}$ of a class $U(\Lambda,d)\alpha\in U(\Lambda,d)\backslash J(A,d)$ the fractional ideal
$I(\alpha):=\bigcap_{\wp\in Spec(R)}\Lambda_{\wp}\cdot\alpha(\wp).$
Likewise we get the completed version. Then, $I(\alpha)\simeq I(\beta)$ as fractional differential graded ideal if and only if there is $x\in A^{\times}\cap\ker(d)$ with $\alpha=\beta x$. If $A$ is semisimple, then we may apply Proposition 5.10 we have that
$I(\alpha)_{\wp}=\Lambda\cdot\alpha(\wp)$
for all $\wp\in Spec(R)$.
In the definition below we shall need to consider $K_{0}((\Lambda,d)-LF-dgmod)$. This is defined as the quotient of isomorphism classes of locally free differential graded $(\Lambda,d)$-modules modulo the relation $[(X,\delta_{X})]-[(Y,\delta_{Y})]-[(Z,\delta_{Z})]$ whenever there is a short exact sequence
$0\longrightarrow(Y,\delta_{Y})\longrightarrow(X,\delta_{X})\longrightarrow(Z,\delta_{Z})\longrightarrow 0$
Note that this setting makes sense for any differential graded $R$-algebra $(\Lambda,d)$ whenever $R$ is a Dedekind domain with field of fractions $K$.
Definition 7.3.
Let $R$ be a Dedekind domain with field of fractions $K$.
  • Let $(\Lambda,d)$ be a differential graded order in the semisimple differential graded $K$-algebra $(A,d)$.
    • The group $\widehat{U}(\Lambda,d)\backslash\widehat{J}(A,d)/(A^{\times}\cap\ker(d))$ is the group of completion differential graded idèles of the dg-order $(\Lambda,d)$.
    • The group $U(\Lambda,d)\backslash J(A,d)/(A^{\times}\cap\ker(d))$ is the group of differential graded idèles of the dg-order $(\Lambda,d)$.
  • If $(\Lambda,d)$ is a differential graded $R$-algebra, then the class group $Cl((\Lambda,d))$ of $(\Lambda,d)$ is the subgroup of $K_{0}((\Lambda,d)-LF-dgmod)$ generated by elements $[L]-[\Lambda]$ for locally free differential graded $(\Lambda,d)$-lattices $(L,\delta)$ of rank $1$.
As in the classical case we get
Theorem 7.4.
Let $R$ be a Dedekind domain with field of fractions $K$. Let $(\Lambda,d)$ be a differential graded order in the semisimple differential graded $K$-algebra $(A,d)$. Then
$[\Lambda\alpha]+[\Lambda\beta]-2[\Lambda]=[\Lambda\alpha\beta]-[\Lambda]$
for any two completion differential graded idèles $\alpha$ and $\beta$. In particular, the differential graded idèle group is isomorphic to the differential graded class group.
Proof. We need to verify that the constructions of idèles in [25, Theorem 8.5.11] do not lead out of $\ker(d)$. Besides this the proof of [25, Theorem 8.5.11] holds in the dg-concept verbatim as in the non dg-concept, until the very last argument. Let us go through the arguments of [25, Theorem 8.5.11]. As in the original proof we may replace $\alpha$ by a version after having multiplied by an $r\in R$ such that $\alpha$ is an integral idèle. Again there is an integer $k$ such that $\alpha_{\wp}^{-1}\in\wp^{-k}\widehat{\Lambda}_{\wp}$ and an integer $t$ and $x\in A^{\times}\cap\ker(d)$ such that $\beta_{\wp}x-1\in\wp^{t}\widehat{\Lambda}_{\wp}$, and again that $\beta x$ is an integral idèle. We observe that $\beta x\in\ker(d)$. Replacing $\beta$ by $\beta x$ is possible since we do not quit $\ker(d)$ by this operation. Again, by the same computation $\alpha_{\wp}\beta_{\wp}\alpha_{\wp}^{-1}\beta_{\wp}^{-1}$ is invertible and in $\ker(d)$ as $\alpha$ and $\beta$ are. By Lemma 7.2 units which are cycles are precisely the units in the ring of cycles. The morphism $f$ in [25, page 310] is easily seen to be a morphism of dg-modules. The fact that $\ker(f)=\Lambda\alpha\beta$ still holds by the same proof. The fact that $\textup{im}(f)=\Lambda$ still holds also in the dg-version by the very same proof. Hence there is a short exact sequence of locally free dg-modules
$0\longrightarrow\Lambda\alpha\beta\longrightarrow\Lambda\alpha\oplus\Lambda\beta\longrightarrow\Lambda\longrightarrow 0$
and $\Lambda\alpha\beta=\Lambda\beta\alpha$. This proves the theorem.
Remark 7.5.
Note that $\Lambda$ is projective, and therefore
$0\longrightarrow\Lambda\alpha\beta\longrightarrow\Lambda\alpha\oplus\Lambda\beta\longrightarrow\Lambda\longrightarrow 0$
splits as $\Lambda$-modules. However, a differential graded module $(M,\delta)$ is projective in the category of dg-modules if and only if $(M,\delta)$ is exact (cf [1]). Indeed the splitting will not commute with the differential if the complex is not exact.
Remark 7.6.
Since we do not factor out the same space as for the classical idèles, in general the dg-class group is not a subgroup of the usual class group. Indeed, if $\alpha\in J(A,d)$ is represented by a $(\alpha_{\wp}\in A_{\wp}^{\times}\cap\ker(d)\;|\;\wp\in Spec(R))$, and if $u=(u_{\wp})_{\wp\in Spec(R)}\in U(\Lambda)$, then in general $u\alpha=(u_{\wp}\alpha_{\wp})_{\wp\in Spec(R)}$ will not be in $J(A,d)$. Indeed,
$d(u_{\wp}\alpha_{\wp})=d(u_{\wp})\alpha_{\wp}\pm u_{\wp}d(\alpha_{\wp})=d(u_{\wp})\alpha_{\wp}.$
and we have seen in Example 3.5 that $d(u_{\wp})\neq 0$ in general.
However, if $d=0$, the concepts coincide and $Cl(\Lambda,0)=Cl(\Lambda)$.
Corollary 7.7.
Let $R$ be a Dedekind domain with field of fractions $K$ and let $(A,d)$ be a differential graded algebra, semisimple as an algebra. Let $(\Lambda,d)$ be a differential graded order in $(A,d)$. Then $Cl(\Lambda,d)\simeq U(\ker(d))\backslash{{J(K\cdot\ker(d))/(K\cdot\ker(d)^{\times})}}$. In particular, if $\ker(d)$ is a semisimple algbra, then $Cl(\Lambda,d)\simeq Cl(\ker(d))$.
Proof. By Theorem 7.4 we have $Cl(\Lambda,d)\simeq U(\Lambda,d)\backslash J(A,d)/{(A^{\times}\cap\ker(d))}$. By Lemma 7.2 we have $U(\Lambda,d)=U(\ker(d))$ and $A^{\times}\cap\ker(d)=\ker(d)^{\times}.$ The first part is now immediate. If $\ker(d)$ is semisimple, then $\ker(d|_{\Lambda})$ is an $R$-order in $\ker(d)$ and [25, Corollary 8.5.12] gives the result.
More generally, we get the following
Proposition 7.8.
Let $R$ be a Dedekind domain with field of fractions $K$. Let $(A,d)$ be a differential graded semisimple finite dimensional $K$-algebra and let $(\Lambda,d)$ be a differential graded $R$-order in $(A,d)$. Then there is a group homomorphism
$Cl(\Lambda,d)\longrightarrow Cl(\Lambda)$
Proof. Indeed, since $A^{\times}\cap\ker(d)\subseteq A^{\times}$ and since $\Lambda^{\times}\cap\ker(d)\subseteq\Lambda^{\times}$, we get a group homomorphism $J(\Lambda,d)\longrightarrow J(\Lambda)$. Further, by the same argument the same embedding shows that $U(\Lambda,d)$ maps to $U(\Lambda)$ and hence we get a group homomorphism $U(\Lambda,d)\backslash J(A,d)\longrightarrow U(\Lambda)\backslash J(A)$. Again, since $(A^{\times}\cap\ker(d))\subseteq A^{\times}$ the same map yields a group homomorphism
$U(\Lambda,d)\backslash J(A,d)/(A^{\times}\cap\ker(d))\longrightarrow U(\Lambda)\backslash J(A)/A^{\times}$
and hence since $Cl(\Lambda,d\simeq U(\Lambda,d)\backslash J(A,d)/(A^{\times}\cap\ker(d))$ and since $Cl(\Lambda)\simeq U(\Lambda)\backslash J(A)/A^{\times}$, we get the statement.
Example 7.9.
Consider the ${\mathbb{Z}}$-order
$\Lambda={\mathbb{Z}}\cdot\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)+p\cdot Mat_{2\times 2}({\mathbb{Z}})$
in $Mat_{2\times 2}({\mathbb{Q}})$. Then, following Example 3.5 this is a differential graded order with the differential $d_{1}$. Further, by [3, §34D] we have that $Cl(\Lambda)$ is of order $2$ whenever $p-1$ is divisible by $4$ (and trivial otherwise). However, $\ker(d_{1})={\mathbb{Z}}[\epsilon]/\epsilon^{2}$, and hence $Cl(\Lambda,d_{1})$ is trivial as is easily seen using the idèle description. Therefore, the map in Proposition 7.8 is not surjective in general.
Since localisation is flat, we get $(H(\Lambda,d))_{\wp}\simeq H(\Lambda_{\wp},d_{\wp})$. For any dg-module $(M,\delta)$ we get $(H(M,\delta))_{\wp}\simeq H(M_{\wp},\delta_{\wp})$. Moreover, if $(M,\delta)$ is a differential graded $(\Lambda,d)$-module, then by Lemma 1.6 we get that $H(M,\delta)$ is a graded $H(\Lambda,d)$-module. Further, if $(M,\delta)$ is a locally free dg $(\Lambda,d)$-module, then $(M_{\wp},\delta_{\wp})$ is a free $(\Lambda_{\wp},d_{\wp})$-module. Hence $H(M,\delta)$ is a graded locally free $H(\Lambda,d)$-module.
If we identify
$Cl(H(\Lambda,d))=\ker\left(K_{0}(H(\Lambda,d)-LF-mod)\longrightarrow K_{0}(H(A,d)-proj)\right)$
taking homology is therefore a group homomorphism
$Cl(\Lambda,d)\stackrel{{\scriptstyle CH_{(\Lambda,d)}}}{{\longrightarrow}}Cl(H(\Lambda,d)).$
Note however that $H(\Lambda,d)$ is not an order in general.
Definition 7.10.
Let $R$ be the ring of integers in a number field $K$. Let $(\Lambda,d)$ be a differential graded order in the semisimple differential graded $K$-algebra $(A,d)$. Then define the homology-isomorphism class group kernel $\ker(CH_{(\Lambda,d)})=:Cl(\Lambda,d)_{hi}$.
Remark 7.11.
Note that for each locally free differential graded $(\Lambda,d)$-module $(L,\delta)$ we have that
$Cl(\Lambda,d)_{hi}+[(L,\delta)]=CH_{(\Lambda,d)}^{-1}(CH_{(\Lambda,d)}([(L,\delta)])$
parameterizes those locally free dg-$(\Lambda,d)$-modules which have the same homology as $(L,\delta)$. Those which are quasi-isomorphic to $(L,\delta)$ do share the homology with $(L,\delta)$. However, being quasi-isomorphic is stronger in general than just having isomorphic homology.
We now consider locally free $H(\Lambda,d)$-modules.
Lemma 7.12.
Let $R$ be a Dedekind domain with field of fractions $K$ and let $B$ be a finitely generated $R$-algebra. For any $B$-module $M$ let
$t(M):=\{x\in M\;|\;\exists r\in R\setminus\{0\}:rx=0\}$
be the torsion submodule of $M$. If $L$ is a locally free $B$-module of rank $n$, then $t(L)\simeq t(B)^{n}$.
Proof. By hypothesis $L_{\wp}\simeq B_{\wp}^{n}$ for all primes $\wp\in Spec(R)$. Since $R$ is Dedekind, $R_{\wp}$ is a principal ideal domain (cf [25, Lemma 7.5.9]). Hence $t(L_{\wp})=t(L)_{\wp}$. Since $R$ is Dedekind, any non zero prime ideal is maximal (cf [25, Lemma 7.5.15]) and hence the primary decomposition (cf [25, Theorem 7.2.5]) of $t(L)$ gives a direct product decomposition $t(L)=\prod_{\wp\in Spec(R)}t(L)_{\wp}$. We hence get $t(L)\simeq t(B)^{n}$.
Corollary 7.13.
Let $R$ be a Dedekind domain with field of fractions $K$, let $(A,d)$ be a finite dimensional and semisimple and differential graded $K$-algebra. Let $(\Lambda,d)$ be a differential graded $R$-order in $(A,d)$. Then
$Cl(H(\Lambda,d))\simeq Cl(H(\Lambda,d)/t(H(\Lambda,d))).$
In particular, if $H(A,d)=0$, then $Cl(\Lambda,d)_{hi}=0$.
Proof. By Lemma 7.12 for any rank $1$ locally free $H(\Lambda,d)$-module $L$ we have $t(L)=t(H(\Lambda,d))$ and therefore
$H(\Lambda,d)/t(H(\Lambda,d))\otimes_{H(\Lambda,d)}-:Cl(H(\Lambda,d))\longrightarrow Cl(H(\Lambda,d)/t(H(\Lambda,d)))$
is an isomorphism. Note that $H(\Lambda,d)/t(H(\Lambda,d))\otimes_{H(\Lambda,d)}-\simeq(R/tR)\otimes_{R}-$.
If now $(A,d)$ is exact, then $H(\Lambda,d)$ is torsion, and therefore locally free $H(\Lambda,d)$-modules are actually free. This proves the lemma.
We shall need a result due to Wehlen. Recall (cf e.g. [11, Chapter 3]) that the Baer lower radical $L(A)$ of an algebra $A$ is the intersection of all prime ideals of $A$. Every element of $L(A)$ is nilpotent and if $A$ is Noetherian, then $L(A)$ is a nilpotent ideal [11, Theorem 3.11].
Theorem 7.14.
[22, Theorem 2.4] Let $R$ be a Prüfer domain and let $A$ be a finitely generated algebra over $R$. Let $L(A)$ be the (Baer) lower radical of $A$. If $B:=A/L(A)$ is separable over $R$, i.e. $B$ is projective as an $B\otimes_{R}B^{op}$-module, then there are idempotents $e_{p}$ and $e_{t}$ of $A$ such that
$A=\left(\begin{array}[]{cc}e_{p}Ae_{p}&e_{t}Ae_{p}\\ e_{p}Ae_{t}&e_{t}Ae_{t}\end{array}\right)$
and $e_{p}Ae_{p}/e_{p}L(A)e_{p}$ is $R$-projective, such that $\left(\begin{array}[]{cc}0&e_{t}Ae_{p}\\ e_{p}Ae_{t}&e_{t}Ae_{t}\end{array}\right)\subseteq t(A)$, the $R$-torsion ideal of $A$.
Note that obviously $e_{p}$ and $e_{t}$ are orthogonal and $e_{p}+e_{t}=1$. Further, note that $A/t(A)=e_{p}Ae_{p}/t(e_{p}Ae_{p})$.
Corollary 7.15.
Suppose that $t(R)=0$ and the hypotheses of Theorem 7.14. Then any unit $u$ in $A/t(A)$ can be lifted to a unit $u_{2}$ of $A$.
Proof. If $u$ is a unit of $A/t(A)$, then we get that $u$ is actually a unit of $e_{p}Ae_{p}/t(e_{p}Ae_{p})$. Further, since $e_{p}Ae_{P}/e_{p}L(A)e_{p}$ is $R$-projective, we have that $t(e_{p}Ae_{p})\subseteq e_{p}L(A)e_{p}$ and since the right hand side of the inclusion is a nil ideal, so is the left hand side. Hence there is $u_{0}$ and $v_{0}$ in $e_{p}Ae_{p}$ such that $u_{0}$ maps to $u$ in $A/t(A)$ and such that $u_{1}v_{1}-1$ is in $e_{p}L(A)e_{p}$, whence nilpotent. But since $1+n$ for a nilpotent element $n$ is a unit, we have that $u_{1}$ is a unit. Therefore we may find a unit $u_{1}\in e_{p}Ae_{p}$ such that $u_{1}$ maps to $u$ in $A/t(A)$. But then there is a unit $u_{2}:=\left(\begin{array}[]{cc}u_{1}&0\\ 0&1\end{array}\right)$ in $A$ (actually in the Pierce decomposition above) such that $u_{2}$ maps to $u$ in $A/t(A)$.
Theorem 7.16.
Let $R$ be a Dedekind domain with field of fractions $K$ and let $(A,d)$ be a differential graded algebra such that $A$ is finite dimensional separable over $K$. Suppose that $\ker(d)$ is separable as a graded algebra. Let $(\Lambda,d)$ be a differential graded $R$-order in $(A,d)$. Denote the canonical map $\pi:\ker(d)\longrightarrow H(\Lambda,d)$ and denote $\overline{H(\Lambda,d)}:=H(\Lambda,d)/t(H(\Lambda,d))$. Then $\overline{H(\Lambda,d)}$ is a classical $R$-order in $H(A,d)$, and we have an exact sequence
$0\longrightarrow Cl(\Lambda,d)_{hi}\longrightarrow Cl(\Lambda,d)\longrightarrow Cl(\overline{H(\Lambda,d)})\longrightarrow 0$
Proof. It is a classical fact that separable algebras are semisimple. Since $\ker(d)$ is a semisimple algebra, Lemma 4.2 shows that we see that $H(A,d)$ is a semisimple algebra. Hence, $H(\Lambda,d)/t(H(\Lambda,d))$ is an $R$-order in the semisimple $K$-algebra $H(A,d)$.
We need to see that the right hand map is surjective. For this we shall use the interpretation of class groups as idèles.
We hence need to see that
$\widehat{U}(\Lambda,d)\backslash\widehat{J}(A,d)/(A^{\times}\cap\ker(d))\longrightarrow\widehat{U}(\overline{H(\Lambda,d)})\backslash\widehat{J}(H(A,d))/H(A,d)^{\times}$
is surjective.
First, by Lemma 4.2 any unit of $H(A,d)$ lifts to a unit in $\ker(d)$. Now, clearly $\ker(d)^{\times}\subseteq(A^{\times}\cap\ker(d))$. For a finitely generated $R$-algebra $B$ satisfying that $B/L(B)$ is separable, if $\overline{u}$ is a unit in $B/tB$, then by Corollary 7.15 any unit $u$ in $B/tB$ lifts to a unit $u$ in $B$. Now, by hypothesis $\ker(d)$ is finite dimensional separable. This is equivalent to the fact that in the Wedderburn decomposition the centres of the skew fields are separable extensions of the base field. Hence $H(A,d)$ is separable as well. Therefore $H(\Lambda_{\wp},d)/t(H(\Lambda_{\wp},d))$ is separable for all $\wp\in Spec(R)$, and so is its quotient modulo its Baer (lower) radical. Following Corollary 7.15 we can hence lift any unit of $H(\Lambda_{\wp},d)/t(H(\Lambda_{\wp},d))$ to a unit of $H(\Lambda_{\wp},d)$
By semisimplicity $A_{\wp}^{\times}\cap\ker(d)\longrightarrow H(A_{\wp},d)^{\times}$ is surjective. By the same argument $\widehat{J}(A,d)\longrightarrow\widehat{J}(H(A,d))$ is surjective as well. Since $\pi(\widehat{U}(\Lambda,d))\subseteq\widehat{U}(H(\Lambda,d))$, we get indeed a map $Cl(\Lambda,d)\longrightarrow Cl(\overline{H(\Lambda,d)})$.
Now, for any locally free $\overline{H(\Lambda,d)}$-ideal $\overline{I}$ we can find an idèle $\overline{\alpha}\in J(\overline{H(A,d)})$ such that $\overline{I}=\bigcap_{\wp\in Spec(R)}\overline{H(\Lambda,d)}\cdot\overline{\alpha}_{\wp}$. Since $\widehat{J}(A,d)\longrightarrow\widehat{J}(H(A,d))$ is surjective, there is an idèle $\alpha\in J(A,d)$ which maps to $\overline{\alpha}$. But then the locally free ideal
$I:=\bigcap_{\wp\in Spec(R)}\Lambda\cdot\alpha_{\wp}$
maps to $\overline{I}$. This shows the surjectivity of $Cl(\Lambda,d)\longrightarrow Cl(\overline{H(\Lambda,d)})$. The statement on the kernel of this map follows by definition.
8. Reducing to dg-orders in differential graded simple algebras
A major tool in studying class groups of orders is the following result.
Theorem 8.1.
(Reiner-Ullom [20])(Mayer-Vietoris theorem for class groups of orders) Let $R$ be a Dedekind domain with field of fractions $K$ and let $\Lambda$ be an $R$-order in the finite dimensional semisimple $K$-algebra $A$ satisfying the Eichler condition. Let $e^{2}=e\in Z(A)$ be a non trivial central idempotent of $A$. Put $f:=1-e$. Then $\Lambda e$ is an $R$-order in $Ae$ and $\Lambda f$ is an $R$-order in $Af$ and we have a pullback diagram
${\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot e}$$\scriptstyle{\cdot f}$${\Lambda\cdot e\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{e}}$${\Lambda f\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{f}}$${\overline{\Lambda}}$
for $\overline{\Lambda}:=\Lambda e/\Lambda\cap\Lambda e\simeq\Lambda f/\Lambda\cap\Lambda f$ and $\pi_{e}$, resp. $\pi_{f}$ being the canonical morphisms. Further, there is a group homomorphism $\delta$ such that $\delta$ and the canonical maps induce an exact sequence
$\Lambda^{\times}\longrightarrow(\Lambda e)^{\times}\times(\Lambda f)^{\times}\longrightarrow\overline{\Lambda}^{\times}\stackrel{{\scriptstyle\delta}}{{\longrightarrow}}Cl(\Lambda)\longrightarrow Cl(\Lambda e)\times Cl(\Lambda f).$
Keep the assumptions and notations of Theorem 8.1 for the moment. In addition if $(A,d)$ is a finite dimensional semisimple differential graded $K$-algebra, and if $(\Lambda,d)$ is a differential graded $R$-order in $(A,d)$, then by Remark 2.7 we have $d(e)=0=d(f)$ and hence $(\Lambda e,d)$ and $(\Lambda f,d)$ are differential graded orders in $(Ae,d)$ respectively $(Af,d)$. Since $\overline{\Lambda}=\Lambda e/(\Lambda\cap\Lambda e)$, and since $d$ induces a differential on both $\Lambda$ and on $\Lambda e$, it also induces a differential $\overline{d}$ on $\overline{\Lambda}$. Therefore, the maps in
${(\Lambda,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot e}$$\scriptstyle{\cdot f}$${(\Lambda\cdot e,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{e}}$${(\Lambda\cdot f,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{f}}$${(\overline{\Lambda},\overline{d})}$
are still well-defined maps of differential graded algebras. Since
${\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot e}$$\scriptstyle{\cdot f}$${\Lambda\cdot e\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{e}}$${\Lambda\cdot f\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{f}}$${\overline{\Lambda}}$
is a pullback diagram of $R$-algebras, the diagram
${(\Lambda,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot e}$$\scriptstyle{\cdot f}$${(\Lambda\cdot e,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{e}}$${(\Lambda\cdot f,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{f}}$${(\overline{\Lambda},\overline{d})}$
is a pullback diagram of differential graded algebras. Recall that $Cl(\Lambda,d)$ is generated by $[L]-[\Lambda]$ for $L$ being rank $1$ locally free differential graded modules. Hence the map
$Cl(\Lambda,d)\stackrel{{\scriptstyle Cl(e)\times Cl(f)}}{{\longrightarrow}}Cl(\Lambda e,d)\times Cl(\Lambda f,d)$
is still well-defined. We recall the definition of $\delta$. Let $u\in\overline{\Lambda}^{\times}$. Then consider the pullback diagram
${L_{u}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{e}}$$\scriptstyle{\alpha_{f}}$${\Lambda\cdot e\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{e}}$${\overline{\Lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot u}$${\Lambda\cdot f\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{f}}$${\overline{\Lambda}}$
and define $\delta(u):=[L_{u}]-[\Lambda]$.
In the dg-case, if $u\in\ker(\overline{d})\cap\overline{\Lambda}^{\times}$, then we get again a pullback diagram
${(L_{u},d_{u})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{e}}$$\scriptstyle{\alpha_{f}}$${(\Lambda\cdot e,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{e}}$${(\overline{\Lambda},\overline{d})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot u}$${(\Lambda\cdot f,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{f}}$${(\overline{\Lambda},d)}$
defining a locally free differential graded $(\Lambda,d)$-module $(L,d_{u})$. We may define a map $\delta_{d}$ by $\delta_{d}(u):=[(L_{u},d_{L})]-[(\Lambda,d)]$ in this case. It remains to show that $\ker(Cl(e)\times Cl(f))=\textup{im}(\delta_{d})$.
Suppose that $u\in\ker(d)\cap\overline{\Lambda}^{\times}$. Then
$Cl(f)\circ\delta_{d}(u)=[L_{u}\cdot f,d_{u}]-[\Lambda f,d)]=[\alpha_{f}(L_{u},d_{u})]-[\Lambda f,d)]=[\Lambda f,d)]-[\Lambda f,d)]=0$
and likewise $Cl(e)\circ\delta_{d}(u)=0$.
Let $(L,d_{L})$ be a rank $1$ locally free dg $(\Lambda,d)$-module in the kernel of $Cl(e)\times Cl(f)$. Let $\alpha$ be the idèle of $(L,d_{L})$. Then we know that $\alpha\cdot e$ is the principal idèle and also $\alpha\cdot f$ is the principal idèle in the corresponding algebras. Now
$A^{\times}\cap\ker d=(Ae^{\times}\cap\ker d)\times(Af^{\times}\cap\ker d).$
Further, $\ker(d)=\ker(de)\times\ker(df)$. We hence only need to see consider ordinary idèles in order to examine the kernel of $Cl(e)\times Cl(f)$. But then, by the classical Mayer-Vietoris-like theorem for class groups of orders we obtain that
$\ker(Cl(e)\times Cl(f))=\textup{im}(\delta_{d}).$
We hence obtained the following
Theorem 8.2.
Let $R$ be a Dedekind domain with field of fractions $K$ and let $(\Lambda,d)$ be a differential graded $R$-order in the finite dimensional semisimple differential graded $K$-algebra $(A,d)$, and suppose that $A$ satisfies the Eichler condition. Let $e^{2}=e\in Z(A)$ be a non trivial central idempotent of $A$. Put $f:=1-e$. Then $(\Lambda e,de)$ is a differential graded $R$-order in $(Ae,de)$ and $(\Lambda f,df)$ is a differential graded $R$-order in $(Af,df)$ and we have a pullback diagram
${\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot e}$$\scriptstyle{\cdot f}$${\Lambda\cdot e\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{e}}$${\Lambda\cdot f\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{f}}$${\overline{\Lambda}_{e}}$
for $\overline{\Lambda}_{e}:=\Lambda e/\Lambda\cap\Lambda e\simeq\Lambda f/\Lambda\cap\Lambda f$ and $\pi_{e}$, resp. $\pi_{f}$ being the canonical morphisms. Further, there is a group homomorphism $\delta_{d}$ such that $\delta_{d}$ and the canonical maps induce an exact sequence
$\Lambda^{\times}\cap\ker(d)\longrightarrow((\Lambda e)^{\times}\times(\Lambda f)^{\times})\cap\ker(d)\longrightarrow\overline{\Lambda}_{e}^{\times}\cap\ker(\overline{d})\stackrel{{\scriptstyle\delta_{d}}}{{\longrightarrow}}Cl(\Lambda,d)\longrightarrow Cl(\Lambda e,de)\times Cl(\Lambda f,df).$
Proof. The result follows from the comments leading to the exactness of
$\overline{\Lambda}_{e}^{\times}\cap\ker(\overline{d})\stackrel{{\scriptstyle\delta_{d}}}{{\longrightarrow}}Cl(\Lambda,d)\longrightarrow Cl(\Lambda e,de)\times Cl(\Lambda f,df).$
Since $\Lambda$ is an order in a semisimple algebra $A$ and $e^{2}=e\in Z(A)$, the sequence
$\Lambda^{\times}\stackrel{{\scriptstyle\gamma}}{{\longrightarrow}}((\Lambda e)^{\times}\times(\Lambda f)^{\times})\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}\overline{\Lambda}_{e}^{\times}$
is exact by the classical situation. Since
${\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot e}$$\scriptstyle{\cdot f}$${\Lambda\cdot e\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{e}}$${\Lambda\cdot f\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{f}}$${\overline{\Lambda}_{e}}$
is a pullback diagram, we get an induced sequence of group homomorphisms
$\Lambda^{\times}\cap\ker(d)\longrightarrow((\Lambda e)^{\times}\times(\Lambda f)^{\times})\cap\ker(d)\stackrel{{\scriptstyle\pi_{d}}}{{\longrightarrow}}\overline{\Lambda}_{e}^{\times}\cap\ker(\overline{d})$
which composes to 00, and where $\pi_{d}$ is the restriction of $\pi$ to $((\Lambda e)^{\times}\times(\Lambda f)^{\times})\cap\ker(d)$. If $x\in\ker(\pi_{d})$, then there is $v\in\Lambda^{\times}$ with $\gamma(v)=x$. We need to see that $v\in\ker(d)$. However, $K\Lambda=K\Lambda e\times K\Lambda f$ and $K\overline{\Lambda}=0$. Hence,
$\textup{id}_{K}\otimes_{R}\gamma=\textup{id}_{A}$
which shows that $x\in\ker(d)$ implies $v\in\ker(d)$. This proves the theorem.
Remark 8.3.
Taking homology we can compare the Mayer-Vietoris sequence of Theorem 8.2 with the Mayer-Vietoris sequence of Theorem 8.1. Denoting as before
$\overline{H(\Lambda,d)}=H(\Lambda,d)/tH(\Lambda,d),$
and likewise for the other orders occurring in the statement, observing that there is a map $\ker(\overline{d})\longrightarrow H(\overline{\Lambda}_{e},\overline{d})$, we get a commutative diagram
[…]
In the situation of Theorem 8.2 we shall need a particular subgroup of the unit group of $\overline{\Lambda}$.
Definition 8.4.
Let $R$ be a Dedekind domain with field of fractions $K$ and let $(\Lambda,d)$ be a differential graded $R$-order in the finite dimensional semisimple differential graded $K$-algebra $(A,d)$. Let $e^{2}=e\in Z(A)$ and $f=1-e$. Let $D_{e}$ be the image of $((\Lambda e)^{\times}\times(\Lambda f)^{\times})\cap\ker(d)\longrightarrow\overline{\Lambda}^{\times}\cap\ker(\overline{d})$. Put
$X_{e}:=\ker((\overline{\Lambda}_{e}^{\times}\cap\ker\overline{d})\rightarrow\overline{H(\overline{\Lambda}_{e},\overline{d})}^{\times})$
we define homology-isomorphism units of $\overline{\Lambda}_{e}:=\Lambda e/(\Lambda\cap\Lambda e)$ as
$\overline{\Lambda}_{hi}^{\times}(e):=X_{e}/\langle[X_{e},X_{e}],D_{e}\rangle.$
Proposition 8.5.
Let $R$ be a Dedekind domain with field of fractions $K$ and let $(\Lambda,d)$ be a differential graded $R$-order in the finite dimensional semisimple differential graded $K$-algebra $(A,d)$. Suppose that $H(A,d)$ is semisimple as a graded algebra. Let $e^{2}=e\in Z(A)$ be a non trivial central idempotent of $A$. Put $f:=1-e$ and $\overline{\Lambda}:=\Lambda e/(\Lambda\cap\Lambda e)$. Then the sequence
$\overline{\Lambda}^{\times}_{hi}(e)\longrightarrow Cl_{hi}(\Lambda,d)\longrightarrow Cl_{hi}(\Lambda e,de)\times Cl_{hi}(\Lambda f,df)$
is exact.
Proof. The proof only uses standard diagram chasing. Here are the details. Recall from Remark 8.3 that there is a commutative diagram with exact rows
[…]
Since the class groups are abelian groups, we can consider the abelianisations of the various unit groups, denoted by the index $ab$, for short. Put $F_{e}$ the image of
$\overline{H(\Lambda e,de)}^{\times}\times\overline{H(\Lambda f,df)}^{\times}\longrightarrow\overline{H(\overline{\Lambda},\overline{d})}^{\times}.$
This then gives a commutative diagram with exact rows (where we replace $D_{e}$ and $F_{e}$ by the image in the abelianisation
[…]
Taking kernels of the vertical maps can be extended to a commutative diagram with exact columns
[…]
Clearly $\beta_{K}\circ\alpha_{K}=0$, which shows $\textup{im}(\alpha_{K})\subseteq\ker(\beta_{K})$.
If $x\in\ker(\beta_{K})$, then
$\iota_{2}(x)=\delta_{d}(y)\in\ker(\epsilon)\cap\ker(\pi_{2})$
for some $y\in\overline{\Lambda}^{\times}\cap\ker(\overline{d})$ and
$(\overline{\delta}\circ\pi_{1})(y)=(\pi_{2}\circ\delta_{d})(y)=\pi_{2}(x)=0.$
Since $\overline{\delta}$ is injective, $\pi_{1}(y)=0$, which shows that there is $y_{1}\in\overline{\Lambda}^{\times}_{hi}$ such that $\iota_{1}(y_{1})=y$. Then
$(\iota_{2}\circ\alpha_{K})(y_{1})=\delta_{d}(\iota_{1}(y_{1}))=\delta_{d}(y)=\iota_{2}(x).$
Since $\iota_{2}$ is injective, $\alpha_{K}(y_{1})=x$. This shows the statement.
Acknowledgements
I wish to thank Bernhard Keller for encouragement in an early stage of the project, and for useful comments on the first draft of the manuscript. I also wish to thank Yann Palu for indicating reference [18] to me.